### v1

# Revisiting Lossless Convexification: Theoretical Guarantees for Discrete-time Optimal Control Problems

Abstract
Lossless Convexification (LCvx) is a modeling approach that transforms a class of nonconvex optimal control problems, where nonconvexity primarily arises from control constraints, into convex problems through convex relaxations. These convex problems can be solved using polynomial-time numerical methods after discretization, which converts the original infinite-dimensional problem into a finite-dimensional one. However, existing LCvx theory is limited to continuous-time optimal control problems, as the equivalence between the relaxed convex problem and the original nonconvex problem holds only in continuous-time. This paper extends LCvx to discrete-time optimal control problems by classifying them into normal and long-horizon cases. For normal cases, after an arbitrarily small perturbation to the system dynamics (recursive equality constraints), applying the existing LCvx method to discrete-time problems results in optimal controls that meet the original nonconvex constraints at all but no more than $n_{x}-1$ temporal grid points, where $n_{x}$ is the state dimension. For long-horizon cases, the existing LCvx method fails, but we resolve this issue by integrating it with a bisection search, leveraging the continuity of the value function from the relaxed convex problem to achieve similar results as in normal cases. This paper improves the theoretical foundation of LCvx, expanding its applicability to real-world discrete-time optimal control problems.
1 Introduction
LCvx is a technique that reformulates a class of nonconvex continuous-time optimal control problems—where nonconvexity primarily stems from control constraints—into convex problems through novel relaxation methods. Pontryagin’s maximum principle provides a theoretical guarantee for this approach, ensuring that the solution to the relaxed convex problem still satisfies the constraints of the original nonconvex problem (Açıkmeşe and Ploen, 2005; Berkovitz, 1975). By converting complex nonconvex problems into convex ones, this relaxation reduces computational cost as the resulting convex problem, after being discretized into a finite-dimensional form, can then be efficiently solved using standard solvers such as ECOS (Domahidi et al., 2013) or Mosek (ApS, 2019) in polynomial time to a specified accuracy (Nemirovski, 2004).
LCvx research has been active for the past decade, and it has proven useful in numerous applications  (Malyuta et al., 2022). The initial application of the LCvx method targeted Mars pinpoint landings (Açıkmeşe and Ploen, 2005, 2007), where the nonzero lower bound on the magnitude of the control input vector led to nonconvex control constraints. Later, the LCvx family has been generalized to handle a wide range of continuous-time optimal control problems, including but not limited to those with generalized nonconvex input constraints (Açıkmeşe and Blackmore, 2011), input pointing constraints (Carson et al., 2011), integer constraints (Harris, 2021; Echigo et al., 2023, 2024), as well as problems with additional convex affine and quadratic state constraints (Harris and Açıkmeşe, 2013, 2014). Moreover, the efficacy of the LCvx method has also been tested in real-world experiments (Açıkmeşe et al., 2012, 2013) and is now being used as a baseline for Moon landing applications(Fritz et al., 2022; Shaffer et al., 2024).
Despite the success of LCvx in real-world applications, its theoretical results have a gap, which may have practical implications for its use. While LCvx theory establishes an equivalence between a continuous-time nonconvex optimal control problem and its convex relaxation, both remain continuous-time problems and cannot be directly solved by numerical solvers, which handle only finite-dimensional problems. To address this, we discretize both problems into finite-dimensional, discrete-time versions (Açıkmeşe and Blackmore, 2011). However, this discretization introduces additional challenges, as LCvx theory does not guarantee that the optimal solution of the discretized convex problem satisfies the constraints of the discretized nonconvex problem.
In fact, the optimal solution of the discrete-time convex relaxed problem can violate the constraints of the original discrete-time nonconvex problem. In this paper, we provide examples where the optimal solution of the relaxed convex problem violates the nonconvex constraints of the original problem at single, multiple, or all temporal grid points. These examples emphasize the necessity of establishing a solid theory for directly applying the LCvx method to discrete-time nonconvex optimal control problems.
Instead of considering a general class of nonconvex control constraints, we focus on a simple problem where the lower bound of the nonzero norm on the control input magnitude is the sole source of nonconvexity. Here, the input magnitude is defined by a convex function of the input variable. We chose this problem because it can demonstrate the essence of LCvx method without introducing excessive mathematical complexity. The application of the LCvx method to a broader class of nonconvex problems will be explored in future studies.
The primary goal of this paper is to develop a new LCvx method that integrates the existing LCvx method with a bisection search. Additionally, we establish a comprehensive proof framework to validate this method for discrete-time nonconvex optimal control problems. In this framework, we classify these problems into two categories: normal cases and long-horizon cases. Long-horizon cases arise when the control horizon is sufficiently long to admit an optimal solution for the convexified problem, but the input magnitude of this solution consistently violates the nonconvex input constraints throughout the trajectory. All other problems are classified as normal cases. For normal cases, our method is identical to the existing LCvx method, as the existing approach has demonstrated effective performance for such cases (Açıkmeşe and Ploen, 2005, 2007). However, in long-horizon cases, the existing approach fails to provide feasible solutions for the nonconvex problem. To address this, we introduce a bisection search strategy specifically designed to handle long-horizon cases.
For the normal case, we provide a theoretical guarantee for applying the existing LCvx method to discrete-time optimal control problems by proving a bound on the number of points where the optimal solution to the convexified problem can violate the original nonconvex constraints. Specifically, we show that, after an arbitrarily small perturbation to the dynamics, the optimal solution to the convexified problem violates nonconvex constraints at no more than $n_{x}-1$ temporal grid points, where $n_{x}$ is the state dimension. Furthermore, we show that the perturbed problem maintains all necessary assumptions from the relaxed convex problem and we establish bounds on the impact of the perturbation. Numerical examples are also provided to demonstrate the necessity of the perturbation and the tightness of the upper bound. Our proof technique is closely related to the discretized Pontryagin’s Maximum Principle (Sethi, 2021; Canon et al., 1970), as both approaches are based on the Lagrangian function for finite-dimensional optimal control problems. However, unlike the discretized Pontryagin’s Maximum Principle, our method does not involve the use of the Hamiltonian.
In contrast, the existing LCvx method does not work for long-horizon cases. To address this, we design a new LCvx method by combining the earlier LCvx method with a bisection search. This approach is motivated by the fact that existing LCvx theory requires strict transversality assumptions ((Malyuta et al., 2022, Condition 2), (Açıkmeşe and Blackmore, 2011, Condition 2)), which fail in long-horizon cases. To overcome this, we divide the control horizon into two phases. In the first phase, we use an arbitrary feasible control input to ensure minimum cost. The final state of the first phase is then used as the initial state for the second phase, where we apply the existing LCvx method to the resulting nonconvex optimal control problem. We expect the control horizon for the second phase to be short enough to satisfy the transversality assumption and belong to normal cases, while still being long enough to maintain minimum input cost. To achieve this balance, we developed a bisection search to find the optimal transition time between phases, based on the fact that the value function of the resulting convex optimal control problem for the entire trajectory is continuous. By employing a bisection search, the second-phase control problem falls into the normal case category, enabling us to directly apply the theoretical results of normal cases to long-horizon problems. We note that another method in the LCvx family also uses the bisection search (Kunhippurayil et al., 2021) for continuous-time systems. The main distinction of our work is that we focus on discrete-time optimal control problems, including terminal costs in the objective function, and impose affine constraints on the terminal state.
Statement of Contributions: In this paper, we develop a new LCvx method for discrete-time nonconvex optimal control problems, bridging the gap between LCvx theory and its real-world applications. Specifically, our contributions are as follows.
First, we develop a new method and proof framework for discrete-time optimal control problems, which differs from previous LCvx research and other asymptotic theories for discretizations of continuous-time optimal control (Ober-Blöbaum et al., 2011; Lee et al., 2022). By focusing on discrete-time settings, we can provide a tight upper bound for the number of temporal grid points at which the optimal control violates the nonconvex constraints after applying the LCvx method.
Second, our proof framework employs linear algebra and finite-dimensional convex analysis, thus avoiding the need for infinite-dimensional mathematical tools such as Pontryagin’s maximum principle. This approach not only simplifies the theoretical development but also offers a fresh perspective on the mechanism of LCvx.
Third, our proof framework eliminates the requirement for the transversality assumption found in earlier LCvx theories ((Malyuta et al., 2022, Condition 2) and (Açıkmeşe and Blackmore, 2011, Condition 2)), an assumption that is often difficult to verify before computation. This removal makes our method more applicable to real-world applications.
Road map: In Section 2, we review the background of LCvx research and define the optimization problem under consideration. Next, in Section 3, we provide an intuitive explanation of the mechanism of LCvx using convex analysis techniques and define the mathematical criteria for normal cases and long-horizon cases. In Section 4, we discuss the validity of LCvx for normal cases, define perturbations to the discrete dynamics, and provide theoretical guarantees for the LCvx method on the perturbed nonconvex optimal control problem. In Section 5, we address the long-horizon case and propose a bisection search algorithm for these cases. Finally, in Section 6, we present numerical experiments to demonstrate counterexamples where the LCvx method is invalid and showcase the perturbation and bisection search algorithms.
1.1 Notation
[…]
2 Background
In this section, we discuss the background of the LCvx method. Consider the optimization problem mentioned in (Malyuta et al., 2022, Problem 6), which includes no state constraints other than boundary conditions. We further simplify this problem to a time-invariant linear system without external disturbances. We believe that this basic form captures the essential ideas of the LCvx method without adding mathematical complexity. This leads to the following optimal control problem:
Problem 1:
[…]
$\operatorname{s.t.}$ $\dot{x}(t)=A_{c}x(t)+B_{c}u(t),$
$\rho_{\min}\leq g_{1}(u(t)),\quad g_{0}(u(t))\leq\rho_{\max},\text{for }t\in[0,t_{f}]$
$x(0)=x_{\text{init}},\quad G(x(t_{f}))=0.$
Here, ${m}\colon\mathbb{R}^{n_{x}}\to\mathbb{R},\ {l_{c}}\colon\mathbb{R}\to\mathbb{R},\ {g_{0}}\colon\mathbb{R}^{n_{u}}\to\mathbb{R}$, and ${g_{1}}\colon\mathbb{R}^{n_{u}}\to\mathbb{R}$ are assumed to be convex $C^{1}$ smooth functions. ${G}\colon\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{G}}$ is an affine map. The state trajectory ${x}\colon[0,t_{f}]\to\mathbb{R}^{n_{x}}$ and the control input trajectory ${u}\colon[0,t_{f}]\to\mathbb{R}^{n_{u}}$ satisfy a continuous ordinary differential equation constraint. The condition $G(x(t_{f}))=0$ represents the boundary condition for the state trajectory $x$. $A_{c}\in\mathbb{R}^{n_{x}\times n_{x}}$ and $B_{c}\in\mathbb{R}^{n_{x}\times n_{u}}$ are constant matrices. $t_{f}\in\mathbb{R}$ is a free flight time and $x_{\text{init}}\in\mathbb{R}^{n_{x}}$ is a fixed initial state. $\rho_{min}\in\mathbb{R}$ and $\rho_{max}\in\mathbb{R}$ are fixed constants. The nonconvexity of Problem 1 arises solely from potentially nonconvex constraints $\rho_{\min}\leq g_{1}(u(t))$. The functions $g_{0}$ and $g_{1}$ characterize the magnitude of the input variable. A classic realization for both is to use the norm function, ensuring that the feasible region of Problem 1 meets the condition $\rho_{\min}\leq\|u(t)\|\leq\rho_{\max}$ for all $t$. These constraints arise from planetary soft landing applications, where the net thrust of rocket engines must exceed a positive lower bound to prevent shutdown (Açıkmeşe and Ploen, 2005).
The LCvx technique introduced in (Açıkmeşe and Blackmore, 2011) considers the following convex relaxation of the Problem 1:
Problem 2:
[…]
$\;\operatorname{s.t.}$ $\dot{x}(t)=A_{c}x(t)+B_{c}u(t),$
$\rho_{\min}\leq\sigma(t),\ g_{0}(u(t))\leq\rho_{\max},\ g_{1}(u(t))\leq\sigma(t),$
$x(0)=x_{\text{init}},\quad G(x(t_{f}))=0,$
where ${\sigma}\colon[0,t_{f}]\to\mathbb{R}$ is a slack variable introduced to convexify Problem 1. Because Problem 2 is a convex relaxation of Problem 1, the optimal value of the former serves as a lower bound for that of the latter. Theorem (Açıkmeşe and Blackmore, 2011, Theorem 2) demonstrates that, under certain assumptions, an optimal solution to Problem 2 satisfies the constraints of Problem 1 almost everywhere, making it an optimal solution to Problem 1 as well.
Now, we consider the discretized, finite-dimensional, versions of Problem 1 and Problem 2. We apply a temporal discretization with zero-order hold for both infinite-dimensional problems to solve them numerically (Açıkmeşe and Ploen, 2005, 2007; Açıkmeşe and Blackmore, 2011). Meanwhile, having a free final time $t_{f}$ can undermine the convexity of the resulting optimization problem. To address this, a commonly used method is to apply a line search for $t_{f}$. By solving the discretized relaxed problem for each $t_{f}$, one can find a $t_{f}$ with the smallest objective value (Açıkmeşe and Ploen, 2007). Hence, in the rest of this paper, we only consider the cases where $t_{f}$ is fixed.
With the following solution variables:
$x=$$\{x_{1},x_{2},\cdots,x_{N+1}\},\ x_{i}\in\mathbb{R}^{n_{x}}\text{for }i\in\left[1,N+1\right]_{\mathbb{N}}$ (3)
$u=$$\{u_{1},u_{2},\cdots,u_{N}\},\ u_{i}\ \in\mathbb{R}^{n_{u}}\text{for }i\in\left[1,N\right]_{\mathbb{N}}$
$\sigma=$$\{\sigma_{1},\sigma_{2},\cdots,\sigma_{N}\},\ \sigma_{i}\in\mathbb{R}\text{ for }i\in\left[1,N\right]_{\mathbb{N}}.$
The discretized version of Problem 1 is:
Problem 3:
$\begin{array}[]{ll}\underset{x,u}{\min}&m\left(x_{N+1}\right)+\sum_{i=1}^{N}l_{i}\left(g_{1}(u_{i})\right)\\ \operatorname{s.t.}&x_{i+1}=A_{d}x_{i}+B_{d}u_{i},\quad i\in\left[1,N\right]_{\mathbb{N}}\\ &g_{0}(u_{i})\leq\rho_{\max},\ g_{1}(u_{i})\geq\rho_{\min},\quad i\in\left[1,N\right]_{\mathbb{N}}\\ &x_{1}=x_{\text{init}},\quad G\left(x_{N+1}\right)=0.\end{array}$ (4)
The sum of ${l_{i}}\colon\mathbb{R}\to\mathbb{R}$ is an approximation for the integral term in the objective function of Problem 1. Matrices $A_{d}\in\mathbb{R}^{n_{x}\times n_{x}}$ and $B_{d}\in\mathbb{R}^{n_{x}\times n_{u}}$ for this discretized problem are given as (Chen, 1984):
[…]
Since we mainly work with the discretized problem, we simplify the notation by using $A$ for $A_{d}$ and $B$ for $B_{d}$. Condition $G(x_{N+1})=0$ is denoted as the boundary condition, $x_{i+1}=Ax_{i}+Bu_{i}$ as the dynamics constraints, and $m(x_{N+1})+\Sigma_{i}l_{i}(\sigma_{i})$ as the object function.
The discretized version of Problem 2 is:
Problem 4:
$\begin{array}[]{ll}\underset{x,u,\sigma}{\min}&m\left(x_{N+1}\right)+\sum l_{i}\left(\sigma_{i}\right)\\ \operatorname{s.t.}&x_{i+1}=Ax_{i}+Bu_{i},\quad i\in\left[1,N\right]_{\mathbb{N}}\\ &\rho_{\min}\leq\sigma_{i},\ g_{0}(u_{i})\leq\rho_{\max},\ g_{1}(u_{i})\leq\sigma_{i},\ i\in\left[1,N\right]_{\mathbb{N}}\\ &x_{1}=x_{\text{init}},\quad G\left(x_{N+1}\right)=0.\end{array}$ (6)
Problem 4 can be understood as the result of applying the LCvx method to Problem 3. Since Problem 4 is a relaxation of Problem 3, its optimal value serves as a lower bound for that of Problem 3.
Motivated by previous LCvx research on continuous problems, a natural question arises: does the optimal solution to Problem 4 satisfy the constraints of Problem 3? However, this relationship does not hold in general. The numerical counterexamples in Section 6 show that the solution to Problem 4 often does not satisfy the constraints of Problem 3 at certain temporal grid points. Therefore, an optimal solution to Problem 4 is not guaranteed to solve Problem 3. As a result, we shifted our focus to a revised question. At how many temporal grid points can the solution to Problem 4 potentially violate the constraints of Problem 3, thereby invalidating LCvx? Here, we define the validity of LCvx at grid point $i$ as:
Definition 1.
The LCvx method is valid at a grid point $i$ if the control input $u_{i}$ satisfies $g_{0}(u_{i})\leq\rho_{\text{max}}$ and $g_{1}(u_{i})\geq\rho_{\text{min}}$.
We will later prove that, with reasonable assumptions on Problem 4 and with arbitrarily small random perturbations to the discrete dynamics, LCvx is invalid at most at $n_{x}-1$ temporal grid points, where $n_{x}$ denotes the dimension of the variable $x$.
Before the theoretical discussion, we provide some definitions and assumptions for the problem 4 required for our discussion.
Assumption 2.
$l_{i}(\cdot)$ is $C^{1}$ smooth, and increases monotonically on $[\rho_{min},\infty)$. $m$ is $C^{1}$ smooth. Set $\{u|g_{1}(u_{i})\leq\rho_{min}\text{ for all }i\}$ is nonempty. Set $\{u|g_{0}(u_{i})\leq\rho_{max}\text{ for all }i\}$ is a nonempty compact set. For a control at temporal grid point $i$, $g_{0}(u_{i})=\rho_{max}$ implies $g_{1}(u_{i})\geq\rho_{min}$.
Note that if $\{u|g_{1}(u_{i})\leq\rho_{min}\text{ for all }i\}$ is empty, LCvx is trivially valid. The last part of Assumption 2 indicates that in Problem 3, $\{u|g_{1}(u_{i})\leq\rho_{min}\text{ for all }i\}\subset\{u|g_{0}(u_{i})\leq\rho_{max}\text{ for all }i\}$. A similar assumption is required in (Açıkmeşe and Blackmore, 2011). For further details, see the discussion under (Açıkmeşe and Blackmore, 2011, Equation (4)).
We also require that the dynamical system in Problem 4 is controllable:
Assumption 3 (Controllability).
For matrices $A$ and $B$ in Problem 4,
$\operatorname{rank}\left(\begin{array}[]{cccc}B&AB&\cdots&A^{n_{x}-1}B\end{array}\right)=n_{x},$ (7)
i.e., the matrix pair $\{A,B\}$ is controllable.
The controllability of the discretized system is closely related to that of the continuous dynamical system. Under mild assumptions, it has been shown that if the continuous dynamical system is controllable, then its discretized counterpart is also controllable (Chen, 1984, Theorem 6.9).
We now introduce some notation for further assumptions. Since the constraints on $(u,\sigma)$ in Problem 4 are identical at each temporal grid point, we define $V:=$
$\{(\bar{u},\bar{\sigma})\in\mathbb{R}^{n_{u}}\times\mathbb{R}\,|\,\rho_{\min}\leq\bar{\sigma},\,g_{0}(\bar{u})\leq\rho_{\max},\,g_{1}(\bar{u})\leq\bar{\sigma}\},$ (8)
so that the constraint in Problem 4 can be expressed as $(u_{i},\sigma_{i})\in V$ for all $i$. The set $V$ is convex as $g_{0}$ and $g_{1}$ are convex functions. Moreover, we define $V(\cdot)$ as
$V(\bar{\sigma}):=\{\bar{u}\in\mathbb{R}^{n_{u}}\,|\,g_{0}(\bar{u})\leq\rho_{\max},\text{ and }g_{1}(\bar{u})\leq\bar{\sigma}\},$ (9)
i.e. $V(\bar{\sigma})$ is a slice of $V$ at a specific value of $\bar{\sigma}\in\mathbb{R}$. The definition of $V(\cdot)$ establishes a sufficient condition for the validity of LCvx for certain $u_{i}$:
Lemma 4.
Under Assumption 2, if a control $\bar{u}$ lies on the boundary of $V(\bar{\sigma})$ with $\bar{\sigma}\geq\rho_{\min}$, $\bar{u}$ satisfies the constraints in Problem 3, i.e. $g_{0}(\bar{u})\leq\rho_{max}$ and $g_{1}(\bar{u})\geq\rho_{min}$.
{proof}
The continuity of $g_{0}$ and $g_{1}$ implies that either $g_{0}(\bar{u})=\rho_{\max}$ or $g_{1}(\bar{u})=\bar{\sigma}$ holds: If neither condition is met, we can find an open neighborhood of $\bar{u}$ where $g_{0}(u)<\rho_{\max}$ and $g_{1}(u)<\bar{\sigma}$ for all $u$ in this neighborhood. This contradicts the requirement of the theorem that $\bar{u}$ is on the boundary of $V(\bar{\sigma})$. Therefore, when $\bar{\sigma}\geq\rho_{\min}$, under Assumption 2, the conditions $g_{0}(\bar{u})=\bar{\sigma}$ or $g_{1}(\bar{u})=\rho_{\max}$ ensure that $g_{0}(\bar{u})\geq\rho_{\min}$. Consequently, $\bar{u}$ satisfies both the convex and nonconvex constraints of Problem 3.∎
To define other assumptions, we need to further simplify the notation for Problem 4. Let $z:=(x,u,\sigma)\in\mathbb{R}^{n_{z}}$ with $n_{z}=(N+1)n_{x}+Nn_{u}+N$. We denote the object function as ${\Phi}\colon\mathbb{R}^{n_{z}}\to\mathbb{R}$ and encapsulate the dynamics and boundary conditions into an affine constraint $H(z)=0$, where ${H}\colon\mathbb{R}^{n_{z}}\to\mathbb{R}^{n_{H}}$ and $n_{H}$ represents the total number of equality constraints. The remaining inequality constraints are written within the convex set $\tilde{V}=\{(x,u,\sigma)\in\mathbb{R}^{n_{z}}\,|\,(u_{i},\sigma_{i})\in V\text{ for all }i\}$. Thus, Problem 4 can be represented as:
Problem 5:
$\min_{z}$ $\Phi\left(z\right)$ (10)
$\operatorname{s.t.}$ $H\left(z\right)=0,\quad z\in\tilde{V}.$
Since $H$ is an affine function, the Jacobian of $H$, denoted as $\nabla H(z)$, is a constant matrix. Therefore, by eliminating the equations in $H$ if necessary, $\nabla H(z)$ becomes a matrix of full rank. Since the dynamics equalities are designed to be full rank, any elimination of redundant constraints occurs only at the boundary condition and does not affect the dynamics equalities. After this elimination, we make the following assumption:
Assumption 5.
For Problem 5, $\nabla H(z)$ has full row rank.
For the convex Problem 5, we further require the Slater condition:
Assumption 6 (Slater’s condition).
For Problem 5, there exists a vector $z$ in the interior of $\tilde{V}$ such that $H(z)=0$
Assumption 6 is relatively modest. When both $g_{0}$ and $g_{1}$ are Euclidean norm functions, it is equivalent to having a control sequence $\|u_{i}\|<\rho_{max}$ such that the trajectory generated by this $u_{i}$ satisfies $G(x_{N+1})=0$. This holds when the optimal control problem is not overly constrained.
In the following discussion of this paper, Assumption 2, 3, 5, and 6 are valid for Problem 4 without further mentioning.
We conclude the background section with a continuity theorem, which is a major tool in our subsequent theoretical discussion. We begin by introducing a lemma which is a direct consequence of  (Clarke et al., 2008, Lemma 3.3.3, Theorem 3.3.4):
Lemma 7.
Consider an optimization problem
Problem 6:
$\min_{y}$ $M\left(y,p\right)$ (11)
$\operatorname{s.t.}$ $F\left(y,p\right)=0,\quad y\in\Omega,$
where $y\in\mathbb{R}^{n_{y}}$, $p\in\mathbb{R}^{n_{p}}$. ${M}\colon\mathbb{R}^{n_{y}+n_{p}}\to\mathbb{R}$ and ${F}\colon\mathbb{R}^{n_{y}+n_{p}}\to\mathbb{R}^{n_{F}}$ are continuous with respect to $(y,p)$, and $\nabla_{y}F$ exists and is continuous with respect to $(y,p)$. Consider a solution $\left(y^{*},p_{0}\right)$ to the equality/constraint system
$F(y^{*},p_{0})=0,\quad y^{*}\in\Omega,$
such that $\eta=0$ is the unique solution to the following equation
$0\in\nabla_{y}F\left(y^{*},p_{0}\right)^{\top}\eta+N_{\Omega}\left(y^{*}\right).$ (12)
There exist $\delta,\xi>0$ such that for any $p\in\mathbb{R}^{n_{p}}$ with $\|p-p_{0}\|<\xi$, there is a point $y\in\Omega$ satisfying $F(y,p)=0$ and
$\|y-y^{*}\|\leq\frac{\|F(y^{*},p)\|}{\delta}.$
We denote the optimum value to Problem 6 as $M^{*}(p)$. Our goal here is to show the continuity of $M^{*}$.
Theorem 8.
Assume that the function $F$ and an optimal solution $(y^{*},p_{0})$ to Problem 6 satisfy all the assumptions in Lemma 7, $\Omega$ is compact, and $M$ is a locally Lipschitz function with respect to $(y,p)$. Then, The optimal value function ${M^{*}}\colon\mathbb{R}^{n_{p}}\to\mathbb{R}$ to Problem 6 is continuous at $p_{0}$.
{proof}
See Appendix 8.1. ∎
3 Mechanism of LCvx
In this section, we provide a sufficient condition for the validity of LCvx and introduce the long-horizon cases and the normal cases for Problem 4.
To start, we define the Lagrangian of Problem 4 as follows:
$L(x,u,\sigma;\eta,\mu_{1},\mu_{2}):=m\left(x_{N+1}\right)+\sum_{i=1}^{N}I_{V}\left(u_{i},\sigma_{i}\right)$ (13)
$+\sum_{i=1}^{N}l_{i}\left(\sigma_{i}\right)+\sum_{i=1}^{N}\eta_{i}^{\top}(-x_{i+1}+Ax_{i}+Bu_{i}+w_{i})$
$+\mu_{1}^{\top}\left(x_{1}-x_{init}\right)+\mu_{2}^{\top}G\left(x_{N+1}\right),$
where $I_{V}$ is the indicator function of set $V=\{(\bar{u},\bar{\sigma})\in\mathbb{R}^{n_{u}}\times\mathbb{R}\,|\,\rho_{\min}\leq\bar{\sigma},\,g_{0}(\bar{u})\leq\rho_{\max},\,\text{and}\,g_{1}(\bar{u})\leq\bar{\sigma}\}$:
$I_{V}(\bar{u},\bar{\sigma})=\left\{\begin{array}[]{cc}0&\text{if }(\bar{u},\bar{\sigma})\in V,\\ +\infty&\text{otherwise},\end{array}\right.$ (14)
and the dual variable $\eta_{i}\in\mathbb{R}^{n_{x}}$ for $i=1,2,\cdots,N,$ $\eta=[\eta_{1}^{\top},\eta_{2}^{\top},\cdots,\eta_{N}^{\top}]^{\top}\in\mathbb{R}^{n_{x}N},$ $\mu_{i}\in\mathbb{R}^{n_{x}},$ and $\mu_{2}\in\mathbb{R}^{n_{G}}.$
Theorem 9.
Suppose Assumption 6 holds, and let $(x^{*},u^{*},\sigma^{*})$ be a solution of Problem 4. Then, there exist dual variables $(\eta^{*},\mu_{1}^{*},\mu_{2}^{*})$ such that
$\left(x^{*},u^{*},\sigma^{*}\right)=\operatorname{arg}\min_{x,u,\sigma}L\left(x,u,\sigma;\eta^{*},\mu_{1}^{*},\mu_{2}^{*}\right).$ (15)
{proof}
This follows directly from (Clarke, 2013, Theorem 9.4, 9.8 and 9.14). ∎ From equation (15), we derive:
$\sigma^{*}=\operatorname{arg}\min_{\sigma}L\left(x^{*},u^{*},\sigma;\eta^{*},\mu_{1}^{*},\mu_{2}^{*}\right),$ (16)
$u^{*}=\operatorname{arg}\min_{u}L\left(x^{*},u,\sigma^{*};\eta^{*},\mu_{1}^{*},\mu_{2}^{*}\right),$ (17)
$x^{*}=\operatorname{arg}\min_{x}L\left(x,u^{*},\sigma^{*};\eta^{*},\mu_{1}^{*},\mu_{2}^{*}\right).$ (18)
Since in the Lagrangian $L$, $u_{i}$ only appears in terms $I_{V}(u_{i},\sigma_{i})$ and $\eta_{i}^{\top}Bu_{i}$, we can write:
$u_{i}^{*}$ $=\arg\min_{u_{i}}I_{V}\left(u_{i},\sigma_{i}^{*}\right)+\eta_{i}^{*\top}Bu_{i}$ (19)
$=\arg\min_{u_{i}}I_{V\left(\sigma_{i}^{*}\right)}\left(u_{i}\right)+\eta_{i}^{*\top}Bu_{i}.$
Equation (19) characterizes the relationship between $u_{i}^{*}$ and $\sigma_{i}^{*}$. With (19) in hand, we discuss the sufficient condition for LCvx to be valid.
Theorem 10 (Sufficient Condition for LCvx).
Under Assumptions 2 and 6, the condition of $-B^{\top}\eta_{i}^{*}$ being nonzero is sufficient for $u_{i}$ to lie on the boundary of $V(\sigma_{i}^{*})$, thus validating LCvx at the grid point $i$ by Lemma 4. Furthermore, $\eta_{i}^{*}$ is given by $\eta_{i}^{*}=A^{\top(N-i)}\eta_{N}^{*}$.
{proof}
We first address the sufficient condition for $u_{i}^{*}$ to lie on the boundary of $V(\sigma_{i}^{*})$. Given that the right-hand side function of equation (19) is convex in $u_{i}$ and $u_{i}^{*}$ is a minimizer of this convex function, it follows that(Borwein and Lewis, 2006, Proposition 2.1.2)
$-B^{\top}\eta_{i}^{*}\in N_{V(\sigma_{i}^{*})}\left(u_{i}^{*}\right),$
where $N_{V(\sigma_{i}^{*})}\left(u_{i}\right)$ denotes the normal cone of the set $V(\sigma_{i}^{*})$ at point $u_{i}^{*}$. The normal cone has nonzero elements only at the boundary points of $V(\sigma_{i}^{*})$ (Bauschke, 2011, Corollary 6.44). Hence, $-B^{\top}\eta_{i}^{*}$ being nonzero implies that $u_{i}$ lies on the boundary of $V(\sigma_{i}^{*})$. In such a case, LCvx is valid at node $i$ according to Lemma 4.
For the evolution of $\eta_{i}$, considering equation (18), we take the partial derivative of $L$ with respect to $x_{i}$ for all $i$, yielding:
$x_{N+1}:$ $\quad\nabla m\left(x_{N+1}\right)+\mu_{2}^{\top}\nabla G\left(x_{N+1}\right)=\eta^{*}_{N},$ (20a)
$x_{i}:$ $\quad\eta_{i}^{*\top}A-\eta_{i-1}^{*\top}=0,$ (20b)
$x_{1}:$ $\quad\mu_{1}^{\top}+\eta_{1}^{*\top}A=0.$ (20c)
Thus, iteratively applying (20b) leads to $\eta^{*}_{i}=A^{\top(N-i)}\eta^{*}_{N}$. The notation $\eta_{N}$ in (20a), instead of $\eta_{N+1}$, is chosen to simplify the notation in subsequent derivations. ∎
Remark 11.
Theorem 10 offers insight into why discrete LCvx is typically valid in applications. We start by designing a slack variable $\sigma$ such that for a fixed $\sigma$, LCvx is valid as long as the control $u$ is on the boundary of the feasible region defined by $\sigma$. Then, if a dual variable $\eta_{i}^{*}$ is not in the left null space of matrix $B$, the corresponding control $u$ lies on the boundary of the feasible region, thus validating LCvx.
However, Theorem 10 does not guarantee the validity of LCvx, as $-B^{\top}\eta_{i}^{*}=0$ can occur for some $i$. An obvious scenario for $\eta_{i}^{*}=0$ is where $\eta_{N}^{*}=0$. The following lemma characterizes this scenario.
Lemma 12.
Under Assumptions 2 and 6, if $\eta_{N}^{*}=0$, then all $\sigma_{i}^{*}=\rho_{\min}$, and $x_{N+1}$ must be an optimal solution to the following optimization problem:
Problem 7:
$\min_{z}$ $m\left(x_{N+1}\right)$ (21)
$\operatorname{s.t.}$ $G\left(x_{N+1}\right)=0.$
{proof}
From Theorem 10, $\eta_{N}^{*}=0$ implies $\eta_{i}^{*}=0$ for all $i$. By Theorem 9, we have
$(\sigma^{*},u^{*})=\operatorname{arg}\min_{\sigma,u}\sum_{i=1}^{N}l_{i}\left(\sigma_{i}\right)+\sum_{i=1}^{N}I_{V}\left(u_{i},\sigma_{i}\right).$
Given Assumption 2, we have that $\sigma_{i}^{*}=\rho_{\min}$ and $u_{i}\in V(\rho_{\min})$.
We now prove the result on $x_{N+1}$. Equation (20a) ensures that
$\nabla m\left(x_{N+1}\right)+\mu_{2}^{\top}\nabla G\left(x_{N+1}\right)=\eta^{*}_{N}=0,$
indicating that $x_{N+1}$ is a KKT point of the convex optimization Problem 7 and thus a minimizer (Borwein and Lewis, 2006). ∎
Theorem 10 provides a sufficient condition for $\eta_{N}^{*}\neq 0$, leading to a natural division of Problem 4 into two scenarios:
Definition 13.
Problem 4 is defined as the long-horizon case if $\sigma_{i}^{*}=\rho_{\min}$ and $x_{N+1}$ solves Problem 7. If these conditions are not satisfied, Problem 4 is classified as the normal case.
The term long-horizon is used because $\eta_{N}^{*}=0$ typically occurs when the control horizon is large. For a problem in the normal case, $\eta_{N}^{*}\neq 0$ according to Lemma 12. In previous LCvx research, $\eta_{N}^{*}\neq 0$ is usually ensured by assumptions, for example (Malyuta et al., 2022, Condition 2) and (Açıkmeşe and Blackmore, 2011, Condition 2). By identifying the long-horizon case, we can omit such types of assumptions.
In the next section, we discuss the theoretical guarantee for normal cases. The long-horizon cases will be left for the subsequent section.
4 LCvx on Normal cases
In this section, we concentrate on problems in the normal case and thus we assume that
Assumption 14.
Problem 4 is normal, i.e., either $\sigma_{i}>\rho_{\min}$ for some $i$ or $x_{N+1}$ is not the minimizer for Problem 7.
As discussed previously, Assumption 14 implies that $\eta_{N}^{*}\neq 0$. A direct consequence of this assumption is:
Corollary 15.
Given Assumptions 2, 6 and 14, if the matrix $A$ and $B$ are invertible, then LCvx is valid at all temporal grid points.
{proof}
When $\eta_{N}^{*}\neq 0$, since $\eta^{*}_{i}=A^{\top(N-i)}\eta^{*}_{N}$, a full rank matrix $A$ ensures that $\eta_{i}^{*}$ is non-zero for all $i$. The invertibility of $B$ implies that $\eta_{i}^{*\top}B$ is non-zero for every $i$, satisfying the conditions of Theorem 10 for LCvx to hold. ∎
However, Corollary 15 might not always be applicable because it requires matrix $B$ to be a square matrix, which means that the control and state dimensions must be equal; In practice, the state dimension often significantly exceeds the control dimension.
When the state dimension is larger than the control dimension, the validity of LCvx at all grid points is not guaranteed. Numerical examples in Section 6 show the invalidity of LCvx at multiple grid points for a non-square matrix $B$, even when the LCvx problem is normal. Therefore, we aim to establish a weaker statement: the original nonconvex constraints are violated at most at $n_{x}-1$ temporal grid points, where $n_{x}$ is the state dimension. However, this statement is still invalid as counterexamples exist (see Section 6.2). Thus, we consider an even weaker statement: after adding a random perturbation to the dynamics matrix $A$ in Problem 4, the nonconvex constraints are violated at no more than $n_{x}-1$ temporal grid points.
4.1 Pertubation
We define an arbitrarily small perturbation in the context of matrix $A$’s Jordan form, where $A$ is defined in the dynamics constraints of Problem 3 and Problem 4:
Definition 16.
Let matrix $A$’s Jordan form be
$A=Q^{-1}JQ.$
Assuming $A$ has $d$ distinct eigenvalues, we let these eigenvalues as $(\lambda_{1},\cdots,\lambda_{d})$. Define $\tilde{A}(q)$ as a perturbation of $A$ with respect to a vector $q=(q_{1},q_{2},\cdots,q_{d})\in\mathbb{R}^{d}$ if:
  • $\bullet$
    $\tilde{A}(q)=Q^{-1}\tilde{J}Q$, where $\tilde{J}$ is a Jordan matrix that is identical to $J$ except for its diagonal elements.
  • $\bullet$
    $\tilde{A}(q)$ has $d$ eigenvalues $(\tilde{\lambda}_{1},\cdots,\tilde{\lambda}_{d})$ such that
    $(\tilde{\lambda}_{1},\cdots,\tilde{\lambda}_{d})=(\lambda_{1},\lambda_{2},\cdots,\lambda_{d})+(q_{1},q_{2},\cdots,q_{d}).$
  • $\bullet$
    The index of Jordan blocks in $\tilde{J}$ corresponding to $\tilde{\lambda}_{i}$ remains the same as in $J$ corresponding to $\lambda_{i}$.
In essence, perturbing a matrix involves altering its eigenvalues. We intentionally preserve the Jordan form as it may convey important structural information about the dynamical system.
We now perturb Problem 4 by replacing the matrix $A$ in the dynamic constraints with $\tilde{A}(q)$, where $\tilde{A}(q)$ represents a perturbation of $A$ with respect to the vector $q\in\mathbb{R}^{d}$. Consequently, the optimization problem can be reformulated as follows:
Problem 8:
$\begin{array}[]{ll}\underset{x,u,\sigma}{\min}&m\left(x_{N+1}\right)+\sum_{i=1}^{N}l_{i}\left(\sigma_{i}\right)\\ \operatorname{s.t.}&x_{i+1}=\tilde{A}(q)x_{i}+Bu_{i},\quad i\in\left[1,N\right]_{\mathbb{N}}\\ &\rho_{\min}\leq\sigma_{i},\ g_{0}(u_{i})\leq\rho_{\max},g_{1}(u_{i})\leq\sigma_{i},\ i\in\left[1,N\right]_{\mathbb{N}}\\ &x_{1}=x_{init},\quad G\left(x_{N+1}\right)=0.\end{array}$ (22)
If $q=0,$ we restore Problem 4.
In the remainder of this paper, $\mathbb{D}\subset\mathbb{R}^{n}$ denotes an $n$-dimensional unit cube, defined as $\mathbb{D}=\{x\in\mathbb{R}^{n}\mid-1\leq x_{i}\leq 1,\forall i=1,\ldots,n\}$, where $x_{i}$ is the $i$-th component of $x$ in $\mathbb{R}^{n}$. For $\epsilon>0$, we define the scaled cube $\epsilon\mathbb{D}$ as $\{x\in\mathbb{R}^{n}\mid-\epsilon\leq x_{i}\leq\epsilon,\forall i=1,\ldots,n\}$. The following theorem ensures that Assumptions 5, 3, 6, and 14 remain valid for the perturbed problem 8 under sufficiently small perturbations $q\in\epsilon\mathbb{D}$.
Lemma 17.
Suppose that Assumptions 2, 3, 5, 6 and 14 hold for Problem 8 when $q=0$. Then there exists an $\epsilon>0$ such that for any $q\in\epsilon\mathbb{D}\subset\mathbb{R}^{d}$, Assumptions 2, 3, 5, 6 and 14 still hold for Problem 8 after perturbation by $q$.
{proof}
See Appendix 8.2. ∎
Remark 18.
In Lemma 17, we choose $q$ from a cube region instead of the more conventional ball region, because we later want to generate $q$ via a uniform distribution, and the cube region ensures independence for each dimension of $q$.
Lemma 17 establishes that Problem 8 follows the same assumptions as Problem 4. Therefore, all discussions in Section 3 regarding Problem 4 are also applicable to Problem 8 for sufficiently small $q$.
4.2 Quantify the LCvx invalidation
The goal for this section is to demonstrate that when the perturbation $q$ is chosen randomly, the probability that LCvx is invalid for problem 8 at no more than $n_{x}-1$ points is zero. To begin, we examine the necessary conditions for Problem 8 to violate the original nonconvex constraints at the grid points indexed by $P_{1},P_{2},\cdots,P_{k}\in\mathbb{N}$.
Lemma 19.
For Problem 8 with a fixed perturbation $q$, we define
$S=\left(\begin{array}[]{llll}\tilde{A}(q)^{N-P_{1}}B&\tilde{A}(q)^{N-P_{2}}B&\cdots&\tilde{A}(q)^{N-P_{k}}B\end{array}\right),$
where $P_{1},P_{2},\cdots,P_{k}\in\mathbb{N}$ are temporal grid points. Suppose that Assumptions 2, 6, and 14 hold for Problem 8. If LCvx is invalid at the temporal grid points $P_{1},P_{2},\cdots,P_{k}\in\mathbb{N}$, then the rank of $S$ satisfies $\operatorname{rank}(S)<n_{x}$.
{proof}
According to Theorem 10, the necessary condition for LCvx to be violated at the $i$-th point is: $\eta_{i}^{\top}B=0$ Given that $\eta_{i}^{\top}=\eta_{N}^{\top}\tilde{A}(q)^{N-i}$, it follows that $\eta_{N}^{\top}$ must lie in the left null space of $\tilde{A}(q)^{N-i}B$. Consequently, if LCvx is invalid at $\{P_{1},P_{2},\cdots,P_{k}\}$, $\eta_{N}^{\top}$ must lie in the left null space of the matrix $S$. According to Assumption 14, $\eta_{N}$ is nonzero. Hence $S$ has nontrivial left null space. Then rank$(S)<n_{x}$. ∎
Now we have the main theorem.
Theorem 20.
Given that Assumptions 2, 5, 6 and 14 hold for Problem 4, along with the controllability of $\{A,B\}$. Then, there exists an $\epsilon>0$ such that if a vector $q\in\mathbb{R}^{d}$ is uniformly distributed over cube $\epsilon\mathbb{D}\in\mathbb{R}^{d}$, the probability of the optimal trajectory for Problem 8 violating the nonconvex constraints in Problem 3 at more than $n_{x}-1$ grid points is zero.
{proof}
See Appendix 8.3. The key to this proof is to use Lemma 19. We show that for a sufficiently small perturbation $q$ and for arbitrary integers $11\leq P_{1}<P_{2}<\cdots<P_{n_{x}}<N+$ 1, where $N$ is the total amount of temporal grid points, the matrix $S$ defined in Lemma 19 is of full rank with probability one.∎
Remark 21.
The perturbation mentioned in Theorem 20 is solely for the purposes of mathematical discussion and is not required in practical applications. However, such a perturbation is mathematically necessary. The numerical example in Section 6.2, artificially constructed as a counterexample, demonstrates that LCvx can become invalid at more than $n_{x}-1$ nodes if no perturbation is applied. Furthermore, this numerical example also indicates that $n_{x}-1$ is a tight upper bound for the number of temporal grid points at which LCvx becomes invalid, given that the control dimension is one.
We also want to estimate the influence of the perturbation. Taking the control sequence that is optimum to Problem 8, and applying it to the unperturbed dynamics, we are expecting that when perturbation $q$ is sufficiently small, the perturbation does not lead to a significant difference in the output of the unperturbed dynamics.
Theorem 22.
Suppose Assumptions 2, 5, 6 and 14 hold for Problem 8. Let the control sequence obtained by solving Problem 8 as $u(q)$ and apply this control to the unperturbed dynamics, leading to a trajectory
$\tilde{x}_{N+1}(q)$ $=x_{1}+\sum_{i=1}^{N}A^{i-1}Bu_{i}(q)$
$=x_{1}+\sum_{i=1}^{N}\tilde{A}^{i-1}(0)Bu_{i}(q),$
where $u_{i}(q)$ is the control sequence $u(q)$ at the $i$-th temporal grid point. Then, for any $\delta>0$, there exists an $\epsilon>0$ such that for any $\|q\|<\epsilon$, $\tilde{x}_{N+1}(q)$ and $u(q)$ satisfy:
  • $\bullet$
    $\|G(\tilde{x}_{N+1}(q))\|\leq\delta$,
  • $\bullet$
    $m(\tilde{x}_{N+1}(q))+\sum_{i=1}^{N}l_{i}(g_{1}(u_{i}(q)))\leq m^{*}_{o}+\delta$,
where $m^{*}_{o}$ is the optimum value of the discretized nonconvex Problem 3.
{proof}
See Appendix 8.4. ∎
5 LCvx on Long-horizon Cases
In this section, we address long-horizon cases. In such scenarios, the optimal control inputs of Problem 4 can violate the nonconvex constraints of Problem 3 at all temporal grid points, i.e., $g_{1}(u_{i})<\rho_{\text{min}}$ for all $i$. In this case, we expect that an optimal solution to the continuous-time nonconvex Problem 1 would have control inputs $u(t)$ satisfying $g_{1}(u)=\rho_{\text{min}}$, and the final state $x(t_{f})$ is the optimal solution for Problem 7.
To find an optimal solution for the continuous-time nonconvex Problem 1, we adopt a strategy that involves dividing the trajectory into two phases, with the switching point between these phases denoted $t_{s}$. The first phase is defined on the interval $[0,t_{s}]$. In this phase, we set the control input $u:[0,t_{s}]\rightarrow\mathbb{R}^{n_{u}}$ to be constant, ensuring that $g_{1}(u)=\rho_{\text{min}}$, while the state trajectory follows the continuous dynamics described in Problem 1. The final state of the first phase then serves as the initial state for the second phase. In the second phase, we formulate a discrete-time convex optimization problem by discretizing the continuous dynamics over $[t_{s},t_{f}]$. The boundary state constraint is maintained, and the convexified control constraints are evaluated at temporal grid points.
Identifying the precise transition time, $t_{s}$, poses a challenge. We aim for $t_{s}$ to be sufficiently long so that the discrete convex optimization problem in the second phase belongs to the normal cases. Simultaneously, $t_{s}$ must be short enough to minimize the control and final state costs of the second phase. Hence, we adopt the bisection search method to determine $t_{s}$. We will demonstrate that the optimal total cost of the first and second phases is a continuous function of $t_{s}$, which justifies the use of the bisection search method. An optimal $t_{s}$ should minimize the total cost while ensuring that the discrete convex LCvx problem in the second phase is normal. In particular, the continuous analog of this bisection search method approach is discussed in (Kunhippurayil et al., 2021).
As explored in Section 4, solving the discrete LCvx problem does not ensure LCvx validation at all temporal grid points. Thus, we aim to find a control sequence such that the constraint $\rho_{\text{min}}+\epsilon\geq g_{1}(u_{i})\geq\rho_{\text{min}}$ for any arbitrarily small fixed $\epsilon>0$ is violated at no more than $n_{x}-1$ grid points after the perturbation described in Theorem 20. Additionally, $x_{N+1}$ should be an optimal solution to Problem 7 up to any arbitrarily small fixed constant $\epsilon$. The mathematical definition for this statement is available in Theorem 27.
Now, we mathematically define the bisection search method. In the first phase, we fix the control input as a constant $u_{s}\in\mathbb{R}^{n_{u}}$ over a duration $t_{s}\in\mathbb{R}$, with the initial condition $x_{s}\in\mathbb{R}^{n_{x}}$ provided in the original problem. The control $u_{s}$ is chosen to ensure $g_{1}(u_{s})=\rho_{\text{min}}$. The continuous dynamics matrices are denoted as $A_{c}$ and $B_{c}$. By plugging in $u_{s}$ into the continuous dynamics, we determine the final state of the first phase, which also serves as the initial condition for the second phase:
[…]
For the second phase, we define a discrete convexified optimization Problem 4, keeping the dynamics discretized as $N$ pieces with zero-order hold control. The discrete dynamics matrices are expressed as:
$A(t_{s})$ $=\exp\left(A_{c}\frac{t_{f}-t_{s}}{N}\right),$ (24)
[…]
where $t_{f}-t_{s}$ is the amount of time remaining for the second phase.
Furthermore, we modify the objective function of Problem 4 by replacing $l_{i}(\sigma_{i})$ with $l(\sigma_{i})\times(t_{f}-t_{s})/N$, where $l$ is chosen such that its derivative on the domain of $\sigma$ is lower-bounded by some $L_{l}>0$. This adjustment simplifies the mathematical derivation. Theorem 27 justifies this modification, ensuring that the quality of the resulting trajectory and control variables remains unaffected.
The optimization problem for the second phase is then formulated as:
Problem 9:
$\underset{x,u,\sigma}{\min}$ $m(x_{N+1})+\sum_{i=1}^{N}l(\sigma_{i})\frac{t_{f}-t_{s}}{N}+t_{s}l(\rho_{\text{min}})$ (26)
$\operatorname{s.t.}\quad$ $x_{i+1}=A(t_{s})x_{i}+B(t_{s})u_{i},\text{ for }i\in\left[1,N\right]_{\mathbb{N}}.$
$\rho_{\min}\leq\sigma_{i},\ g_{0}(u_{i})\leq\rho_{\max},\ g_{1}(u_{i})\leq\sigma_{i},$
for $\text{for }i\in\left[1,N\right]_{\mathbb{N}}.$
$x_{1}=x_{\text{init}}(t_{s}),\quad G(x_{N+1})=0.$
Here, we introduce a constant term $t_{s}l(\rho_{\text{min}})$ to the objective function of the aforementioned optimization problem to account for the cost of the first phase. This addition simplifies the proof while leaving the optimal solution of the second phase unaffected. The value of the optimal objective function, denoted by $v(t_{s})$, depends on $t_{s}$ and represents the total cost of the flight, where the minimum value of $v(t_{s})$ is $m^{*}+t_{f}l(\rho_{\text{min}})$, with $m^{*}$ being the optimal value of Problem 7. Our primary objective is to determine a $t_{s}$ such that $v(t_{s})<m^{*}+t_{f}l(\rho_{\text{min}})+\epsilon$, for a given and arbitrarily small $\epsilon>0$. Additionally, to ensure the applicability of the theorems from the previous section, we require that Problem 9 satisfies the normal case condition.
To simplify the notation for Problem 9, we define $z=(x,u,\sigma)$, the object function as $M(z,t_{s})$, and all affine equalities as $F(z,t_{s})=0$. The remaining inequality constraints are identical to the set $\tilde{V}$ in (4). Thus, our problem can be represented as:
Problem 10:
$\min_{z}$ $M\left(z,t_{s}\right)$ (27)
$\operatorname{s.t.}$ $F\left(z,t_{s}\right)=0,\quad z\in\tilde{V}.$
The optimal value is a function of $t_{s}$, denoted as $v(t_{s})$.
We require the following assumption for Problem 10.
Assumption 23.
Problem 4 belongs to the long-horizon case. Meanwhile, there exists a $t_{b}$ such that $v(t_{b})>m^{*}+t_{f}l(\rho_{\text{min}})$. Assumptions 2, 3, 5, and 6 are valid for Problem 9 generated by $t_{s}\in[0,t_{b}]$.
Lemma 24.
Assumption 23 ensures that Problem 9 generated by $t_{b}$ has a non-zero $\eta_{N}^{*}$.
{proof}
The condition $v(t_{b})>m^{*}+t_{f}l(\rho_{\text{min}})$ ensures that either $x_{N+1}$ is not the solution to Problem 7, or at least one $\sigma_{i}^{*}$ exceeds $\rho_{\text{min}}$. In both cases, $\eta_{N}^{*}\neq 0$ follows by the contrapositive of Lemma 12.∎
Consequently, we have the following continuity result:
Lemma 25.
Assumption 23 ensures that $v(t)$ is continuous over $(0,t_{b}]$.
{proof}
Assumption 23 guarantees that, for each $t$, solutions exist for Problem 10. For each $t$, we select one solution and denote it as $z^{*}(t)$. Consequently, we can apply Theorem 8 to Problem 10 at the point $(z^{*}(t),t)$ for all $t$, with the prerequisites of Theorem 8 being met by Assumption 23 and Lemma 29 in the Appendix. The compactness in Theorem 8 is satisfied by the same method used in Lemma 31. ∎
Since $v(t)$ is a continuous function of $t$, we can apply the bisection search method to find the transition point. Under Assumption 23, $v(0)=m^{*}+t_{f}l(\rho_{\text{min}}),$ and $v(t_{b})>m^{*}+t_{f}l(\rho_{\text{min}}),$ where $m^{*}$ is the optimal value of Problem 7. We let $t_{\text{low}}=0$ and $t_{\text{high}}=t_{b}$. In each iteration, we choose $t_{\text{mid}}=(t_{\text{low}}+t_{\text{high}})/2$. If $v(t_{\text{mid}})>m^{*}+t_{f}l(\rho_{\text{min}}),$ we replace $t_{\text{high}}$ with $t_{\text{mid}}.$ Otherwise, we replace $t_{\text{low}}$ with $t_{\text{mid}}.$ The following theorem is a direct consequence of a standard bisection search method argument, the fact that $v(t)$ is continuous and Lemma 24.
Theorem 26.
Under Assumption 23, for any $\epsilon>0$, the bisection search method finds a $t_{s}^{*}\in(0,t_{b})$ ensuring the following inequality within finite iterations:
$m^{*}+t_{f}l(\rho_{\text{min}})<v(t_{s}^{*})<m^{*}+t_{f}l(\rho_{\text{min}})+\frac{\epsilon}{2},$ (28)
where $m^{*}$ is the optimal value of Problem 7. Meanwhile, the Problem 10 generated by $t_{s}^{*}$ is normal.
Since the Problem 10 generated by $t_{s}^{*}$ is normal, we can now add the perturbation and apply the results from Theorem 20 and 22. As tuning the switch point changes the horizon and discretization of Problem 10, we cannot directly compare the solution of Problem 10 to that of the discrete Problem 3. Therefore, we directly compare the result of Problem 10 to that of the continuous-time Problem 1. Note that the best possible solution for Problem 1 is the one with control $g_{1}(u(t))=\rho_{\text{min}}$ for all $t$ and $m(x(t_{f}))$ equal to $m^{*}$, where $m^{*}$ is the optimal value of Problem 7. The next theorem shows that the solution of Problem 10 with perturbation achieves $m(x(t_{f}))$ almost equal to $m^{*}$, and $g_{1}(u_{i})$ less than $\rho_{\text{min}}$ up to an arbitrarily small $\epsilon>0.$
To do this, we need a few notations consistent with those of Theorem 20. Consider the Problem 9 generated from $t_{s}^{*}$. We add a perturbation $q\in\mathbb{R}^{d}$ to the dynamics matrix $A(t_{s}^{*})$ and denote the perturbated matrix as $\tilde{A}(t_{s}^{*},q)$. The optimal solution of the resulting perturbed optimization problem is denoted as $(x(q),u(q),\sigma(q))$. If we plug in $u(q)$ into the original dynamics, we have
$\tilde{x}_{N+1}(q)$ $=x_{1}+\sum_{i=1}^{N}A\left(t_{s}^{*}\right)^{i-1}B\left(t_{s}^{*}\right)u_{i}(q)$ (29)
$=x_{1}+\sum_{i=1}^{N}\tilde{A}^{i-1}\left(t_{s}^{*},0\right)B\left(t_{s}^{*}\right)u_{i}(q)$
Here, $\tilde{x}_{N+1}(q)$ is denoted to be the endpoint of the propagated trajectory if we apply the control acquired from the perturbed problem to the original dynamics.
Theorem 27.
Assume Assumption 23 is valid, Choose a function ${l}\colon\mathbb{R}\to\mathbb{R}$ for Problem 9 such that the derivative of $l$ is larger than some $L_{l}>0$ within the domain of $\sigma$. Then, for any $\epsilon>0$, we can acquire $t_{s}^{*}\in\mathbb{R}$ and $\delta>0$ such that for any perturbation $q\in\mathbb{R}^{d}$ with $\|q\|<\delta$ applied to Problem 10 generated by $t_{s}^{*}$, the control $u(q)$ and the propagated trajectory $\tilde{x}_{N+1}(q)$ satisfies:
  • $\bullet$
    $\|G(\tilde{x}_{N+1}(q))\|\leq\delta$,
  • $\bullet$
    $g_{1}(u_{i})\geq\rho_{\text{min}}$ for all $i$ up to $n_{x}-1$ points with probability one.
  • $\bullet$
    $\rho_{\text{min}}+\epsilon>g_{1}(u_{i})$ for all $i$,
  • $\bullet$
    $m^{*}+\epsilon>m(\tilde{x}_{N+1}(q))$,
where $m^{*}$ is the solution to Problem 7.
{proof}
See Appendix 8.5
Theorem 27 guarantees the validity of adding perturbation to the dynamics. Consequently, we introduce Algorithm 1, which is capable of handling both normal and long-horizon cases. This algorithm provides a control sequence that at most violates the original nonconvex constraints at $n_{x}-1$ grid points. In practical applications, the perturbation step is likely to remain deactivated since it serves primarily for theoretical completeness and is activated only in artificial edge cases.
Algorithm 1 LCvx Method with Bisection for the Long Horizon Cases
0:  $A_{c}$, $B_{c}$,$A$,$B$, $u_{s}$, $x_{s}$, $\rho_{\text{min}}$, $\epsilon_{t}>0$, $\epsilon_{q}>0$
0:  Optimized time $t_{s}$, state $x$, control $u$, and slack variables $\sigma$
1:  Set $t_{\text{low}}=0$, $t_{\text{high}}=t_{f}$, $t_{s}^{*}=0$, $x_{\text{init}}=x_{s}$
2:  Solve Problem 4 to obtain $x$, $u$, $\sigma$
3:  if $\sigma_{i}>\rho_{\text{min}}$ for some $i$ or $x_{N+1}$ doesn’t solve Problem 7 then
4:     if the number of nodes that LCvx being invalid $\leq n_{x}-1$  then
5:        return  $t_{s}^{*}$, $x$, $u$, $\sigma$
6:     else
7:        Perturb the optimization Problem 4 with randomly generated $\|q\|\leq\epsilon_{q}$
8:        Solve the perturbated Problem 4 to obtain $x$, $u$, $\sigma$
9:        return  $t_{s}^{*}$, $x$, $u$, $\sigma$
10:     end if
11:  end if
12:  Modify the object function $l_{i}$ to $l(\sigma_{i})\times(t_{f}-t_{s})/N$ that satisfies assumption in Theorem 27
13:  Set $t_{s}=(t_{\text{low}}+t_{\text{high}})/2$
14:  while $(t_{\text{high}}-t_{\text{low}})>\epsilon_{t}$ do
15:     Solve Problem 9 defined by $t_{s}$ and $u_{s}$ to obtain $x$, $u$, $\sigma$
16:     if $\sigma_{i}>\rho_{\text{min}}$ for some $i$ or $x_{N+1}$ doesn’t solve Problem 7 then
17:        Set $t_{\text{low}},t_{\text{high}}=t_{\text{low}},t_{s}$
18:     else
19:        Set $t_{\text{low}},t_{\text{high}}=t_{s},t_{\text{high}}$
20:     end if
21:     $t_{s}=(t_{\text{low}}+t_{\text{high}})/2$
22:  end while
23:  $t_{s}^{*}=t_{\text{high}}$
24:  Solve Problem 9 defined by $t_{s}^{*}$ to obtain $x$, $u$, $\sigma$.
25:  if the number of nodes that LCvx being invalid $\leq n_{x}-1$  then
26:     return  $t_{s}^{*}$, $x$, $u$, $\sigma$
27:  else
28:     Perturb the optimization Problem 9 by $\|q\|\leq\epsilon_{q}$ to obtain $x,\ u,\ \sigma$
29:     return  $t_{s}^{*}$, $x$, $u$, $\sigma$
30:  end if
6 Numerical Examples
In this section, we consider three illustrative numerical examples.
The first example involves a well-conditioned dynamics model that satisfies all assumptions outlined in Sections 3 and 4. The numerical experiment demonstrates that LCvx is invalid at one temporal grid point, indicating that the solution to Problem 4 may not necessarily be the solution of Problem 3.
The second example explores an artificial case, which, while not practical, is useful for illustration. In this case, LCvx is initially invalid at $n_{x}$ points. Introducing a perturbation reduces this violation to $n_{x}-1$ points, highlighting the necessity of a perturbation and the tightness of the $n_{x}-1$ bound in Theorem 20.
The third example addresses a long-horizon case. It serves to demonstrate a typical long-horizon scenario and the application of the bisection search method algorithm.
One can show that Assumptions 2, 3, 5, 6 and 14 are satisfied for the first and second example. Testing Slater’s Condition in Assumption 23 directly for the third example is challenging as it requires verification for all $t_{s}\in[0,t_{b}].$ Developing a method to test such an assumption will be left for future research. Other conditions in Assumption 23 are satisfied for the third example.
For our numerical examples, we used Matlab and selected Mosek (ApS, 2019) as our convex solver, due to its high accuracy in solving convex optimization problems. The accuracy of the convex solver is crucial in demonstrating the effectiveness of the perturbation strategy and the bisection search method.
6.1 Well-Conditioned Example
We begin our first example by considering an optimization problem with a four-dimensional state and a two-dimensional control. We choose $N=10$ temporal grids, resulting in eleven state vectors $\{x_{1},x_{2},x_{3},\cdots,x_{11}\}$ and ten control vectors $\{u_{1},u_{2},\cdots,u_{10}\}$. We use the notation $x_{i,j}$ to denote the $j$-th dimension of the vector $x_{i}$. In the continuous dynamics, we have $t_{f}=10$ and
$A_{c}=\left(\begin{array}[]{cccc}0_{2\times 2}&I_{2\times 2}\\ 0_{2\times 2}&0_{2\times 2}\\ \end{array}\right),\quad B_{c}=\left(0_{2\times 2},\ I_{2\times 2}\right)^{\top},$ (30)
thus the discretized dynamics are given by
$A=I_{4\times 4}+A_{c}\quad B=\left(0.5I_{2\times 2},\ I_{2\times 2}\right)^{\top}.$ (31)
The object function is $x_{11,3}^{2}+x_{11,4}^{2}+\sum_{i=1}^{11}\sigma_{i}$. The functions $g_{1}$ and $g_{2}$ are the squared $L_{2}$ norms; $\rho_{\text{min}}=1$ and $\rho_{\text{max}}=2$; $x_{\text{init}}=[1,2,3,4]^{\top}$ and the boundary condition is $(x_{11,1},x_{11,2})=[1,1]^{\top}$. In summary, we consider:
$\min_{x,u,\sigma}\quad x_{11,3}^{2}+x_{11,4}^{2}+\sum_{i=1}^{10}\sigma_{i}$ (32)
$\operatorname{s.t.}\quad x_{i+1}=Ax_{i}+Bu_{i},\quad i\in\left[1,10\right]_{\mathbb{N}}$
$1\leq\sigma_{i},\quad\|u_{i}\|^{2}\leq\sigma_{i},\quad\|u_{i}\|^{2}\leq 2,\quad i\in\left[1,10\right]_{\mathbb{N}}$
$x_{1}=[1,2,3,4]^{\top},\quad(x_{11,1},x_{11,2})=[1,1]^{\top}$
The solution trajectory and the magnitude of the controls are available in Figure 1. We observe that LCvx is invalid at the eighth temporal grid. Hence, the solution to Problem 4 does not always satisfy the constraints in Problem 3, even for a well-conditioned problem.
(a) State Trajectory
(b) Control Magnitude
Figure 1: State trajectory and control magnitude for a well-conditioned four-dimensional system over ten temporal grid points. At the eighth grid point, the control norm falls below the minimum threshold, indicating an invalid point for LCvx.
6.2 Artificial normal example
In this section, we consider a three-dimensional problem with $N=10$, $A=\operatorname{diag}(1.2,-2.2,1)$, and $B=(0.4,0.3,0.2)^{\top}.$ Notice that $A$ and $B$ forms a controllable matrix pair.
$\min_{x,u,\sigma}\quad x_{10,3}+\sum_{i=1}^{11}\sigma_{i}$ (33)
$\operatorname{s.t.}\quad x_{i+1}=Ax_{i}+Bu_{i-1},\quad i\in\left[1,10\right]_{\mathbb{N}}$
$x_{\text{init}}=[0,0,0]^{\top},\ x_{11,1}=0.5,\quad x_{11,2}=1$
$1\leq\sigma_{i},\quad\|u_{i}\|^{2}\leq\sigma_{i},\quad\|u_{i}\|^{2}\leq 2,\quad i\in\left[1,10\right]_{\mathbb{N}}$
Here, the dynamics do not come from any continuous system, as $A$ has negative eigenvalues. This counterexample is formed by making the matrix $\Lambda$ in the proof of Theorem 20 singular. The numerical tolerance for Mosek is set to $10^{-15}$. The solution trajectory is available in Figure 2(a), and the magnitude of the control is shown in Figure 2(b). We observe that LCvx is invalid at three points. According to Theorem 20, adding a perturbation to the dynamics reduces the violation of nonconvex constraints to at most two points. We consider the following dynamics with $A_{p}=\operatorname{diag}(10^{-7},0,0)$:
$x_{i+1}=(A+A_{p})x_{i}+Bu_{i-1},\ i\in\left[1,10\right]_{\mathbb{N}}$
which introduces a small change to one element in the dynamics matrix. We maintain the numerical tolerance at $10^{-15}$. The updated trajectory and control sequence are available in Figure 2(c) and Figure 2(d), respectively. As expected, LCvx is invalid at two points after perturbation.
(a) State before perturbation
(b) Control magnitude before perturbation
(c) State after perturbation
(d) Control after perturbation
Figure 2: System dynamics and control response before and after perturbation for an artificial three-dimensional problem with $N=10$ temporal grid points. The top figures display the state trajectory and control magnitude prior to perturbation, revealing LCvx invalid at three points. The bottom figures show the system’s response to a perturbation designed as per Theorem 20, which reduces the number of nonconvex constraint violations to two. The state variable remains essentially unchanged.
6.3 Long-horizon example
Our final example is a long-horizon case. It follows the same setup as in example (32), with the distinction that we have $N=50$ and $t_{f}=50$. The trajectory and control magnitude are available in Figures 3(a) and 3(b). These figures show that $x_{N,3}=x_{N,4}=0$ and all control magnitudes are less than $\rho_{\text{min}}=1$, confirming that this is indeed a long-horizon case.
Now, we apply the bisection search method algorithm 1, setting $\epsilon_{t}=10^{-5}$ and the initial control as $[1,0]^{\top}$. The algorithm terminates with $t_{\text{high}}=9.86$. This suggests maintaining the initial control for 9.86 units of time before starting a control sequence generated by (5). The resulting trajectory and control magnitude are shown in Figures 3(c) and 3(d). The final state is $x_{N+1}=(1,1,10^{-8},3\times 10^{-9})$, with all $\sigma_{i}$ equal to 1 up to the computational error tolerance. The control magnitude plot reveals one LCvx invalidation.
(a) State before bisection search method
(b) Control magnitude before bisection search
(c) State after bisection search method
(d) Control magnitude after bisection search
Figure 3: State trajectory and control magnitude before and after applying the bisection search method in a long-horizon optimization scenario with $N=50$ and $t_{f}=50$. The initial figures indicate a long-horizon state where the control magnitudes are below the threshold $\rho_{\text{min}}=1$. The latter figures demonstrate the effectiveness of the bisection search method algorithm 1, resulting in a trajectory that LCvx violation at one point. The information for $\sigma$ is not available after the perturbation.
7 Conclusion
In this paper, we develop a new Lossless Convexification (LCvx) method and provide theoretical guarantees for its application to discrete-time nonconvex optimal control problems, addressing a gap in the existing LCvx literature. Specifically, we demonstrate that under mild assumptions, the optimal solution of the relaxed convex problem satisfies the constraints of the original nonconvex problem at all but at most $n_{x}-1$ temporal grid points when an arbitrarily small perturbation is introduced into the dynamic constraints, where $n_{x}$ is the state dimension. We also introduce a bisection search method to ensure that our new LCvx method is applicable in long-horizon cases, where the existing LCvx method results in an infeasible control sequence. Additionally, the numerical examples illustrate the necessity of the perturbation and the tightness of the $n_{x}-1$ upper bound, while also demonstrating the effectiveness of the bisection search method.
For future work, we aim to investigate the theoretical guarantees of applying the LCvx method to various forms of discrete-time nonconvex optimal control problems. Furthermore, we plan to extend our results to develop a lossless convexification method that ensures continuous-time constraint satisfaction (Elango et al., 2024).
8 Appendix
8.1 Proof of theorem 8
{proof}
Let $y^{*}$ be the optimal solution to Problem 6 at $p_{0}$, satisfying assumptions in Lemma 7. The existence of $y^{*}$ is a consequence of the compactness of $\Omega$ and the continuity of the object function. Lemma 7 ensures the existence of $\delta>0$ and $\xi>0$ such that for any $p\in\mathbb{R}^{n_{p}}$ with $|p-p_{0}|<\xi$, there is a point $y\in\Omega,F(y,p)=0$ and
$\|y-y^{*}\|\leq\frac{\|F(y^{*},p)\|}{\delta}.$ (34)
Hence, the feasible set for Problem 6 is nonempty for $p$ sufficiently close to $p_{0}.$ Additionally, because $\Omega$ is a compact set, optimal solutions exist for Problem 6 for $p$ sufficiently close to $p_{0}.$
We first prove that $\operatorname*{lim\,inf}_{s\rightarrow p_{0}}M^{*}(s)\geq M^{*}(p_{0})$. Consider a sequence $s_{k}\in\mathbb{R}^{n_{p}}$ sufficiently close to $p_{0}$ such that $\lim_{k}M^{*}(s_{k})=\operatorname*{lim\,inf}_{s\rightarrow p_{0}}M^{*}(s)$ and $\lim_{k}s_{k}=p_{0}$. We denote the optimal solution of the Problem 6 generated by $s_{k}$ as $y_{k}$. Due to the compactness of $\Omega$, there exist a convergent subsequence $y_{k_{n}}$ and a variable $\bar{y}$ such that $\lim_{n}y_{k_{n}}=\bar{y}$ and $\operatorname*{lim\,inf}_{s\rightarrow p_{0}}M^{*}(s)=\lim_{n}M^{*}(s_{k_{n}})=\lim_{n}M(y_{k_{n}},s_{k_{n}})$. Due to the continuity of $F(y,p)$ and the closedness of $\Omega$, $\bar{y}$ is feasible for Problem 6 generated by $p_{0}$. Thus, $\lim_{n}M^{*}(s_{k_{n}})=\lim_{n}M(y_{k_{n}},s_{k_{n}})=M(\bar{y},p_{0})\geq M^{*}(p_{0})$ by the continuity of $M(\cdot,\cdot)$.
Next, we aim to show that $\operatorname*{lim\,sup}_{w\rightarrow p_{0}}M^{*}(w)\leq M^{*}(p_{0})$. As the object function $M(y,p)$ is locally Lipschitz and $y$ lies in a compact set, there exists a Lipschitz constant $L_{M}$ for all feasible $(y,p)$ when $\|p-p_{0}\|<\xi$.
Consider a sequence $w_{k}$ such that $\lim_{k}M^{*}(w_{k})=\operatorname*{lim\,sup}_{w\rightarrow p_{0}}M^{*}(w)$ with $|w_{k}-p_{0}|<\xi$ and $\lim_{k}w_{k}=p_{0}$. Lemma 7 ensures the existence of a corresponding sequence $y_{k}\in\Omega$ that $F(y_{k},w_{k})=0$,
$\|y_{k}-y^{*}\|\leq\frac{\|F(y^{*},w_{k})\|}{\delta},$
and $M^{*}(w_{k})\leq M(y_{k},w_{k})$. Therefore,
$\operatorname*{lim\,sup}_{w\rightarrow p_{0}}$ $M^{*}(w)=\lim_{k}M^{*}(w_{k})$ (35)
$\leq\operatorname*{lim\,sup}_{k}M(y_{k},w_{k})$ (36)
$\leq\operatorname*{lim\,sup}_{k}\left(M(y^{*},w_{k})+L_{M}\|y_{k}-y^{*}\|\right)$ (37)
$\leq M(y^{*},p_{0})+L_{M}\operatorname*{lim\,sup}_{k}\frac{\|F(y^{*},w_{k})\|}{\delta}$ (38)
$=M(y^{*},p_{0})+\frac{\|F(y^{*},p_{0})\|}{\delta}=M^{*}(p_{0}).$ (39)
Here, inequality (37) is due to the Lipschitz property of $M$ and inequality (39) is due to the continuity of $F$.
In summary, we have
$\operatorname*{lim\,sup}_{p\rightarrow p_{0}}M^{*}(p)\leq M^{*}(p_{0})\leq\operatorname*{lim\,inf}_{p\rightarrow p_{0}}M^{*}(p),$
thus establishing the continuity of $M^{*}$. ∎
8.2 Proof for Lemma 17
Since Assumption 2 trivially holds, we start by showing the validity of Assumptions 3 and 5.
Lemma 28.
Suppose that Assumptions 3 and 5 hold for Problem 8 at $q=0$. Then, there exists an $\epsilon>0$ such that for any $q\in\mathbb{R}^{d}$ within the cube $\epsilon\mathbb{D}\in\mathbb{R}^{d}$, where $\mathbb{D}\subset\mathbb{R}^{d}$ is a unit cube, both assumptions are satisfied for Problem 8 generated by $q$.
{proof}
The proof of both assumptions relies on the fact that the determinant of a matrix is a continuous function of elements in that matrix. We demonstrate Assumption 3 for Problem 8, while the nonsingularity of $\nabla_{z}H$ in Assumption 5 follows similarly and thus is omitted.
Define
$S(q):=\left(\begin{array}[]{llll}B&\tilde{A}(q)B&\cdots&\tilde{A}(q)^{n_{x}-1}B\end{array}\right).$ (41)
We aim to show that $S(q)$ retains full rank for all $q$ within $\epsilon\mathbb{D}$. $S(q)$ has full rank if and only if the determinant of $S(q)S(q)^{\top}$ is nonzero. As the elements in matrix $\tilde{A}$ are continuous functions of $q$, the elements and, consequently, the determinant of $S(q)S(q)^{\top}$ are continuous functions of $q$. Given that $\tilde{A}(0)=A$, the controllability of the matrix pair $\{A,B\}$ ensures that $S(0)$ has full rank and that the determinant of $S(0)S(0)^{\top}$ is nonzero. Through continuity, an $\epsilon>0$ exists such that for any $q$ within $\epsilon\mathbb{D}$, the determinant of $S(q)S(q)^{\top}$ remains nonzero, and $S(q)$ maintains full rank. The existence of $\epsilon$ is a corollary of (Rudin, 1964, Theorem 4.8). ∎
[…]
$\min_{z}$ $\Phi(z)$ (42)
s.t. $\tilde{H}(z,q)=0,\quad z\in\tilde{V},$
When $q=0$, we have $\tilde{H}(z,0)=H(z)$, thereby retrieving the original Problem 5.
Since we need lemma 7 and theorem 8 for our discussion, we must demonstrate that their assumptions are valid.
Lemma 29.
Assumption 5 and Assumption 6 for Problem 4 ensures that for any vector $z_{0}$ such that
$\tilde{H}(z_{0},0)=0,\quad z_{0}\in\tilde{V},$
we have that $y=0$ is the unique solution of the following equation:
$0\in\nabla_{z}\tilde{H}\left(z_{0},0\right)^{\top}y+N_{\tilde{V}}\left(z_{0}\right)$ (43)
{proof}
We prove this statement by contradiction. Assume there exists $y\neq 0$ such that
$0\in\nabla_{z}\tilde{H}\left(z_{0},0\right)^{\top}y+N_{\tilde{V}}\left(z_{0}\right).$
Consider the function $g(z)=y^{\top}\tilde{H}(z,0)+I_{\tilde{V}}(z)$, where $I_{\tilde{V}}(z)$ is the indicator function for the set $\tilde{V}$. Thus we have $0\in\partial g(z_{0})$ and $\tilde{H}(z_{0},0)=0$. Since $g(z)$ is a convex function, the minimum of $g(z)$ is zero, achieved at $z_{0}$.
Given Assumption 6, there exists an interior point $z^{\prime}$ in $\tilde{V}$ where $\tilde{H}(z^{\prime},0)=0$ and $N_{\tilde{V}}(z^{\prime})=\{0\}$. Therefore, $z^{\prime}$ is also a minimizer of the function $g$, and thus $0\in\partial g(z^{\prime})=y^{\top}\nabla_{z}\tilde{H}(z^{\prime},0)$. Meanwhile, $\nabla_{z}\tilde{H}(z^{\prime},0)=\nabla_{z}\tilde{H}(z_{0},0)$ because $H$ is an affine function of variable $z$. Consequently, the vector $y$ lies in the left null space of $\nabla_{z}\tilde{H}(z,0)$, which contradicts Assumption 5. ∎
Now, we are ready to prove that Assumption 6 holds after perturbation:
Lemma 30.
Suppose that Assumption 5 and 6 hold for Problem 8 when $q=0$. Then, there exists an $\epsilon>0$ such that for any $q\in\mathbb{R}^{d}$ within the cube $\epsilon\mathbb{D}\in\mathbb{R}^{d}$, where $\mathbb{D}\in\mathbb{R}^{d}$ is an unit cube, such that the perturbed Problem 8 with respect to $q$ also satisfies Assumption 6.
{proof}
Slater’s assumption implies the existence of a $z_{0}$ such that $H(z_{0})=\tilde{H}(z_{0},0)=0$, while $z_{0}$ lies in the interior of $\tilde{V}$. Since Lemma 29 ensures the applicability of Lemma 7, there exist $\epsilon_{1}>0$ and $\delta>0$ such that for all $\|q\|<\epsilon_{1}$, there exists a $z(q)$, determined by $q$, satisfying
$\|z(q)-z_{0}\|\leq\frac{\|\tilde{H}(z_{0},q)\|}{\delta},$
and $\tilde{H}(z(q),q)=0$, thus meeting the dynamics and boundary conditions of Problem 8. Given that $\tilde{H}$ is a continuous function, there exists an $\epsilon>0$, such that when $\|q\|<\epsilon$, $\|z(q)-z_{0}\|$ is sufficiently small for $z(q)$ to lie in the interior of $\tilde{V}$, thereby validating Slater’s condition for the perturbed problem. ∎
Now we show that the perturbed problem is normal.
Lemma 31.
Suppose that Assumptions 2, 5, 6 and 14 hold for Problem 8 and $q=0$. Then there exists an $\epsilon>0$ such that for any $q$ within the cube $\epsilon\mathbb{D}$, Assumption 14 is still valid for Problem 8 after perturbation by $q$.
{proof}
We apply Theorem 8 to show that the optimum value of Problem 8 is continuous with respect to $q$. To satisfy the theorem’s assumptions, we introduce a sufficiently large trivial upper bound on $\sigma$, ensuring the compactness of $\tilde{V}$. Assumption 2 guarantees that the optimal solution remains unchanged with or without this upper bound. Consequently, Assumption 2 and Lemma 29 confirm the validity of assumptions to Theorem 8.
Since when $q=0$ Problem 4 is normal, the optimal value of Problem 4 is larger than $m^{*}+\sum_{i=1}^{N}l_{i}(\rho_{\min})$, where $m^{*}$ is the minimum value of Problem 7. Thus, the continuity of the optimal value function ensures the existence of an $\epsilon>0$ such that for all $q\in\epsilon\mathbb{D}$, the optimal value of Problem 8 generated by $\{\tilde{A}(q),B\}$ is greater than $m^{*}+\sum_{i=1}^{N}l_{i}(\rho_{\min})$. Assumption 2 ensures that $l_{i}$ are monotone, indicating that the perturbed problem is still normal. ∎
8.3 Proof for Theorem 20
The following is a useful linear algebra theorem from (Chen, 1984, Theorem 3.5) that aids us in proving Theorem 20.
Theorem 32.
Consider a function $f(\lambda)$ and a matrix $A\in\mathbb{R}^{n\times n}$. The characteristic polynomial of $A$ is
$\Delta(\lambda)=\operatorname{det}(I-\lambda A)=\prod_{i=1}^{d}(\lambda-\lambda_{i})^{m_{i}},$
where $m_{i}$ are the algebraic multiplicity of eigenvalue $\lambda_{i}$, $n=\sum_{i=1}^{d}m_{i}$, and $d$ is the total number of distinct eigenvalues.
There exists a polynomial $h(\lambda)$ as
$h(\lambda):=v_{0}+v_{1}\lambda+\cdots+v_{n-1}\lambda^{n-1},$ (44)
such that
$f(A)=h(A).$
The unknown coefficients of (44) are found by solving $n$ equations:
$f^{(l)}(\lambda_{i})=h^{(l)}(\lambda_{i})\text{ for }l\in\left[0,m_{i}-1\right]_{\mathbb{N}},i\in\left[1,d\right]_{\mathbb{N}},$ (45)
where
$f^{(l)}(\lambda_{i}):=\left.\frac{d^{l}f(\lambda)}{d\lambda^{l}}\right|_{\lambda=\lambda_{i}},$ (46)
and $h^{(l)}(\lambda_{i})$ is defined in the same manner.
Now we start the proof for 20. {proof}[Proof for Theorem 20 ] If $n_{x}>N$, the statement holds trivially; thus, we assume $n_{x}\leq N$. We choose $\epsilon>0$ sufficiently small that Lemma 17 is valid. Thus, the perturbated Problem 8 has controllable dynamics, is normal, and satisfies Slater’s condition.
Consider the matrix
$S(q):=\left(\begin{array}[]{llll}\tilde{A}(q)^{P_{1}}B&\tilde{A}(q)^{P_{2}}B&\cdots&\tilde{A}(q)^{P_{n_{x}}}B\end{array}\right),$ (47)
for specific integers $11\leq P_{1}<P_{2}<\cdots<P_{n_{x}}<N+$ 1, where $N$ is the total number of grid points in Problem 8, $\tilde{A}(q)$ is the perturbation of $A$ with respect to $q$ and $n_{x}$ is the state dimension.
According to Lemma 19, if $S(q)$ is of full rank, then the Problem 8 generated by $\{\tilde{A}(q),B\}$ satisfies LCvx at least at one point among the set $\left\{N-P_{1},N-P_{2},\cdots,N-P_{n_{x}}\right\}$.
We aim to demonstrate that $S(q)$ is of full rank with probability one, regardless of the choice of $P_{1}$ to $P_{n_{x}}$.
According to Theorem 32, all powers of the matrix $\tilde{A}(q)$ can be written as polynomials of $\tilde{A}(q)$ up to the power $n_{x}-1$, where $n_{x}$ is the dimension of the state vector $x$ and of $A$. Then, there exists a matrix $V(q)\in\mathbb{R}^{n_{x}\times n_{x}}$ such that
$\tilde{A}^{P_{i}}(q)=\sum_{j=1}^{n_{x}}V_{ij}(q)\tilde{A}^{j-1}(q),\quad i\in[1,n_{x}]_{\mathbb{N}}$ (48)
Here, $V_{ij}$ represents the element in the $i$-th row and $j$-th column of $V(q)$.
Now we show that $V(q)$ being invertible is sufficient for $S(q)$ to be of full rank: Let $W(q)$ be the inverse of $V(q),$ Then
$\tilde{A}^{i}(q)=\sum_{j=1}^{nx}W_{ij}\tilde{A}^{P_{i}}(q),\quad i\in[1,n_{x}]_{\mathbb{N}}$ (49)
If $S(q)$ has a nonzero left null vector $\xi$, we have $\xi\tilde{A}(q)^{P_{i}}B=0\text{ for all }$ for all $i\in\left[1,n_{x}\right]_{\mathbb{N}}.$ Equation (49) implies that $\xi\tilde{A}^{i}(q)B=0,$ for all $i\left[0,n_{x}-1\right]_{\mathbb{N}}$ contradicting controllability ensured by Lemma 17.
Therefore, $V(q)$ being invertible is sufficient for $S(q)$ to be of full rank. We thus have
$\operatorname{Prob}(\operatorname{rank}(S(q))<n_{x})\leq\operatorname{Prob}(V(q)\text{ being singular})$
The rest of this proof is to show that the probability on the right-hand side is zero.
Setting the distinct eigenvalues of $A$ as $\lambda_{1},\lambda_{2},\cdots,\lambda_{d}$, the eigenvalues of the perturbed matrix $\tilde{A}(q)$ have eigenvalues $\tilde{\lambda}_{i}=\lambda_{i}+q_{i}$ for $i=\{1,2,\cdots,d\}$. Here, $d$ is the number of distinct eigenvalues of $A$.
We start by exploring how $V(q)$ is determined by the eigenvalues $\tilde{\lambda}_{i}=\lambda_{i}+q_{i}$. The characteristic polynomial of $A$ can be written as: $\prod_{i=1}^{d}(\lambda-\lambda_{i})^{m_{i}}$ where $\sum_{i=1}^{d}m_{i}=n_{x}$ with $m_{i}$ being the algebraic multiplicity of $\lambda_{i}$.
Due to the perturbation preserving the Jordan form, the characteristic polynomial of $\tilde{A}$ becomes:
$\prod_{i=1}^{d}(\lambda-\tilde{\lambda}_{i})^{m_{i}}=\prod_{i=1}^{d}(\lambda-\lambda_{i}-q_{i})^{m_{i}},$
where $\sum_{i=1}^{d}m_{i}=n_{x}.$ It is clear that the probability of
$\tilde{\lambda}_{i}=\lambda_{i}+q_{i}=\lambda_{j}+q_{j}=\tilde{\lambda}_{j}$
holding for any $i$ and $j$ is zero. Consequently, we retain $d$ distinct eigenvalues with probability one.
According to Theorem 32, $V(q)$ consists of the coefficients in (44) for each $P_{i}$. For the $P_{i}$ power of $\tilde{A}(q)$ and an eigenvalue $\tilde{\lambda}_{j}$, the equation (45) for the $i$-th row of $V(q)$ is:
$\quad\quad\tilde{\lambda}_{j}^{P_{i}}=V_{i1}+V_{i2}\tilde{\lambda}_{j}+\cdots+V_{in_{x}}\tilde{\lambda}_{j}^{n_{x}-1},$ (50)
$P_{i}\tilde{\lambda}_{j}^{P_{i}-1}=0+V_{i2}+\cdots+V_{in_{x}}(n_{x}-1)\tilde{\lambda}_{j}^{n_{x}-2},$
[…]
$\prod_{k=0}^{m_{j}-2}(P_{i}-k)\tilde{\lambda}_{j}^{P_{i}-m_{j}+1}=0+\cdots+V_{in_{x}}\prod_{k=1}^{m_{j}-1}(n_{x}-k).$
Combining all equations for different $P_{i}$ and $\tilde{\lambda}_{j}$, there are in total $n_{x}^{2}$ unknown variables in $V(q)$ and $n_{x}^{2}$ equations. We now write all equations into a matrix equality.
For the $P_{i}$-th power and the $j$-th eigenvalue, to simplify the notation, we define $\Lambda_{ij}$ as
$\left[\tilde{\lambda}_{j}^{P_{i}},\,P_{i}\tilde{\lambda}_{j}^{P_{i}-1},\,\cdots,\,\prod_{k=0}^{m_{j}-2}(P_{i}-k)\tilde{\lambda}_{j}^{P_{i}-m_{j}+1}\right],$ (51)
$T_{j}$ as
$\left[\begin{array}[]{cccc}1&0&\cdots&0\\ \tilde{\lambda}_{j}&1&\cdots&0\\ \tilde{\lambda}_{j}^{2}&2\tilde{\lambda}_{j}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ \tilde{\lambda}_{j}^{n_{x}-1}&(n_{x}-1)\tilde{\lambda}_{j}^{n_{x}-2}&\cdots&\prod_{k=1}^{m_{j}-1}(n_{x}-k)\tilde{\lambda}_{j}^{n_{x}-m_{j}}\end{array}\right],$ (52)
and
$V_{i}=\left[V_{i1},V_{i2},\cdots,V_{in_{x}}\right].$ (53)
Thus, we have $\Lambda_{ij}=V_{i}T_{j}.$ Stacking $\Lambda$ and $T$ into matrices, we have
$\Lambda=\left[\begin{array}[]{cccc}\Lambda_{11}&\Lambda_{12}&\cdots&\Lambda_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \Lambda_{n_{x}1}&\Lambda_{n_{x}2}&\cdots&\Lambda_{n_{x}d}\end{array}\right]$ (54)
and
$T=\left[T_{1},T_{2},\cdots,T_{d}\right].$ (55)
According to Theorem 32, the matrix $V(q)$ must satisfy
$\Lambda=VT,$
where $\Lambda$, $V$, and $T$ are $n_{x}$-by-$n_{x}$ matrices and functions of $q$. The determinant of $\Lambda$ is the product of the determinants of $V$ and $T$. Hence, for $V$ to be singular, $\Lambda$ must also be singular.
We establish the non-singularity of $\Lambda$ through induction. To begin, we construct a sequence of matrices $C_{k}\in\mathbb{R}^{M_{k}\times M_{k}}$, for $k=1,2,\cdots,d$ and $M_{k}=\sum_{i=1}^{k}m_{i}$. Each $C_{k}$ is designed to be a square matrix formed from the top-left corner of the matrix $\Lambda$. Specifically $C_{k}$ is defined as:
$C_{k}=\left[\begin{array}[]{ccc}\Lambda_{11}&\cdots&\Lambda_{1k}\\ &\vdots&\\ \Lambda_{M_{k}1}&\cdots&\Lambda_{M_{k}k}\end{array}\right]_{M_{k}\times M_{k}}.$ (56)
As $n_{x}=\sum_{i=1}^{d}m_{i}$, where $d$ is the number of distinct eigenvalues that $A$ has, it follows that $C_{d}=\Lambda$.
We initiate induction by proving that the probability of $C_{1}$ being singular is zero. Substituting the expression of $\Lambda_{ij}$ into the definition of $C_{1}$, we obtain:
$C_{1}=\left[\Lambda_{11}^{\top},\Lambda_{21}^{\top},\cdots,\Lambda_{M_{1}1}^{\top}\right]^{\top}=$ (57)
[…]
The determinant of $C_{1}$ is computed as a sum of signed products of matrix entries. Each summand in this sum represents a product of distinct entries, with no two entries sharing the same row or column. Consequently, all summands contribute to the same power order of $\tilde{\lambda}_{1}$. Therefore, the determinant of $C_{1}$ can be expressed as:
$\operatorname{det}(C_{1})=c_{1}\tilde{\lambda}_{1}^{\sum_{r=1}^{m_{1}}P_{r}-\sum_{r=1}^{m_{1}}(r-1)},$
with constant $c_{1}$.
The expression for $c_{1}$ can be found by setting $\tilde{\lambda}_{1}=1$, as this expression for $C_{1}$ holds for all realizations of $\tilde{\lambda}_{1}$. We denote $E_{1}$ as the matrix that generates $c_{1}$, that is, $c_{1}=\operatorname{det}(E_{1})$. Each element in $E_{1}$ is a polynomial in $P_{i}$, where $i\in\{1,2,\cdots,{m_{1}}\}$. Notably, the polynomials in the same column share the same coefficients. This structure allows $E_{1}$ to be transformed into a Vandermonde matrix through column transformations. Given that $P_{i}$ are distinct, this Vandermonde matrix is nonsingular, implying $c_{1}\neq 0$. Thus $C_{1}$ is nonsingular as long as $\tilde{\lambda}_{1}\neq 0$. Given that $\tilde{\lambda}_{1}$ follows a uniform distribution, the probability of $\tilde{\lambda}_{1}=0$ is zero. Therefore, $C_{1}$ is nonsingular with probability one.
Now, assuming that $C_{k}$ is nonsingular with probability one, we show that $C_{k+1}$ is also nonsingular with probability one. We start by showing that with fixed values of $\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{k}$ and the condition that $C_{k}$ being nonsingular, the conditional probability of $C_{k+1}$ being nonsingular is one. Partition $C_{k+1}$ into the following blocks:
$\left[\begin{array}[]{ll}C_{k}&D_{k+1}\\ R_{k+1}&S_{k+1}\end{array}\right].$ (58)
Here, $R_{k+1}\in\mathbb{R}^{m_{k+1}\times M_{k}}$ does not contain $\tilde{\lambda}_{k+1}$, and $D_{k+1}\in\mathbb{R}^{m_{k+1}\times M_{k}}$ contains terms involving $\tilde{\lambda}_{k+1}$ up to order $P_{M_{k}}$. The terms of $\tilde{\lambda}_{k+1}$ with power higher than $P_{M_{k}}$ are all included in $S_{k+1}\in\mathbb{R}^{m_{k+1}\times m_{k+1}}$.
Given the fixed values of $\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{k}$, the determinant $\operatorname{det}(C_{k+1})$ becomes a polynomial in $\tilde{\lambda}_{k+1}$ with fixed coefficients. The term with the highest power in this polynomial arises from the product of elements in $C_{k}$ with elements in $S_{k+1}$. The absolute value of the highest power term is in fact $|\operatorname{det}(C_{k})\operatorname{det}(S_{k+1})|.$
Note that the matrix $S_{k+1}$ shares the same pattern as $C_{1}$. Following a similar line of reasoning as in the case of $E_{1}$, we can conclude that the coefficient of $\operatorname{det}(S_{k+1})$ is also nonzero. Since $C_{k}$ is nonsingular, the determinant $\operatorname{det}(C_{k+1})$ has a nonzero leading term, implying that it has a finite number of roots. Meanwhile, the distribution of $\tilde{\lambda}_{k+1}$ is a uniform distribution independent of previous eigenvalues. Therefore, the conditional probability of $\tilde{\lambda}_{k+1}$ coinciding with these roots i.e.
$\operatorname{Prob}(C_{k+1}\text{ singular}|C_{k}\text{ nonsingular},(\tilde{\lambda}_{1},\cdots,\tilde{\lambda}_{k})\text{ fixed})$ (59)
is zero. Note that $\operatorname{Prob}(\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{k}|C_{k}\text{ nonsingular})$ is a uniform distribution over the domain of $\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{k},$ which is the set $\epsilon\mathbb{D}$ projected to the first $k$ dimensional minus a measure zero set representing perturbations that make $C_{k}$ singular. By integrating into $\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{k}$, we deduce that:
$\operatorname{Prob}(C_{k+1}\text{ singular}|C_{k}\text{ nonsingular})=0.$
Thus $\operatorname{Prob}(C_{k+1}\text{ singular})=0,$ as $c_{k}$ is nonsigular with probability one.
Using induction, $\Lambda=C_{d}$ is nonsingular with probability one, which implies that $V$ is also nonsingular. Consequently, $S(q)$ has full rank with probability one. Due to the discussion at the start of this proof, the probability of LCvx being invalid at all grid points $(N-P_{1},N-P_{2},\ldots,N-P_{n_{x}})$ is zero. This means that, for any selection of $n_{x}$ grid points, the probability of simultaneously violating the original nonconvex constraints at these points is zero. Hence, the probability of at least $n_{x}$ violations of the original nonconvex constraints is also zero.∎
8.4 Proof for Theorem 22
{proof}
Since $\tilde{A}^{i}(\cdot)$ is a smooth function of the perturbation and the matrix norm is a Lipschitz function, there exists an $\epsilon_{1}$ such that $\|\tilde{A}^{i}(\cdot)\|$ is a Lipschitz function for perturbations $\|q\|\leq\epsilon_{1}$ and for $i=0,1,2,\ldots,N-1$. Let $L_{A}$ denote the maximum Lipschitz constant among all $i$ of $\|\tilde{A}^{i}(\cdot)\|$. Additionally, Assumption 2 ensures that $u_{i}(q)$ is bounded, i.e., $\max\|u_{i}(q)\|<D$ for some $D>0$. Therefore, when $\|q\|\leq\epsilon_{1}$, we can compare $\tilde{x}_{N+1}(q)$ with $x_{N+1}(q)$, and obtain:
$\|\tilde{x}_{N+1}(q)-x_{N+1}(q)\|$
$=\|\sum_{i=1}^{N}(\tilde{A}^{i-1}(0)-\tilde{A}^{i-1}(q))Bu_{i}(q)\|$
$\leq\sum_{i=1}^{N}L_{A}\|q\|\|\|B\|\|u_{i}(q)\|\leq NDL_{A}\|B\|\|q\|.$
Hence, $\tilde{x}_{N+1}(q)$ locates in a local neighborhood of $x_{N+1}(q)$ with radius $NDL_{A}\|B\|\epsilon_{1}$. Meanwhile, since $G$ is an affine function and $G(x_{N+1}(0))=0$, we have $\|G(\tilde{x}_{N+1}(q))\|<\delta$, for suffciently small $\|q\|$.
Now we move on to the second result. Since the control is bounded, $x(q)$ is bounded. Thus, $m(\cdot)$ in Problem 8 is a Lipschitz function within the domain of $x$ because $m$ is $C^{1}$ smooth. with Lipschitz constant $L_{m}$. Thus:
$|m(\tilde{x}_{N+1}(q))-m(x_{N+1}(q))|\leq NDL_{A}L_{m}\|B\|\|q\|.$
Meanwhile, Lemma 29, Lemma 31 (for the compactness), and Theorem 8 indicate that the optimum value of Problem 8 is continuous with respect to $q$. Therefore, for any $\delta>0$ there exists $\epsilon>0$ such that for all $\|q\|<\epsilon,$
$m(\tilde{x}_{N+1}(q))+\sum_{i=1}^{N}l_{i}(g_{1}(u_{i}(q)))$
$\leq m(x_{N+1}(q))+\sum_{i=1}^{N}l_{i}(g_{1}(u_{i}(q)))+K_{q}\|q\|$
$\leq m(x_{N+1}(0))+\sum_{i=1}^{N}l_{i}(g_{1}(u_{i}(0)))+K_{q}\|q\|+\delta/2$
$\leq m^{*}_{o}+K_{q}\|q\|+\delta/2,$
where $K_{q}=NDL_{A}L_{m}\|B\|.$ Here, the second inequality arises from Theorem 8 that the optimum value of Problem 8 is continuous with respect to $q$ , and that $(u(0),x_{N+1}(0))$ is the solution to the unperturbed problem. The last inequality is derived from the fact that the optimum value of Problem 4 lower bounds that of Problem 3. Hence, for sufficiently small $\|q\|$, we have $m(\tilde{x}_{N+1}(q))+\sum_{i=1}^{N}l_{i}(g_{1}(u_{i}(q)))\leq m^{*}_{o}+\delta.$
8.5 Proof for Theorem 27
{proof}
Denote $(x(q),u(q),\sigma(q))$ the solution to Problem 10 generated by $t_{s}^{*}$ and perturbed by $q$. For any $\epsilon>0$, we set $\eta=\min\{\epsilon L_{l}(t_{f}-t_{s}^{*})/N,\frac{3\epsilon}{4}\}$. According to Theorem 26, there exists a $t_{s}^{*}$ such that
$v\left(t_{s}^{*}\right)=$ $m\left(x_{N+1}(0)\right)+\frac{t_{f}-t_{s}^{*}}{N}\sum_{i=1}^{N}l\left(\sigma_{i}(0)\right)$ (60)
$+t_{s}^{*}l\left(\rho_{\min}\right)\leq m^{*}+t_{f}l\left(\rho_{\min}\right)+\frac{\eta}{2}$
where $m^{*}$ is the optimum of Problem 7 and $x_{N+1}(0)$ is the result of the unperturbed Problem 10 formed by $t_{s}^{*}$.
The first result follows from Theorem 22. The second result is from Theorem 20 as long as we choose $\delta$ sufficiently small such that $\|q\|\leq\delta$ is included in the cube area defined by Theorem 20. We now proceed to prove the third and fourth results.
According to Lemma 31,
$m\left(x_{N+1}(q)\right)+\frac{t_{f}-t_{s}^{*}}{N}\sum_{i=1}^{N}l\left(\sigma_{i}(q)\right)+t_{s}^{*}l\left(\rho_{\text{min}}\right)$ (61)
is a continuous function of $q$ in a neighborhood of zero. Therefore, there exists a $\delta>0$ such that for all $\|q\|\leq\delta$,
$m\left(x_{N+1}(q)\right)+\frac{t_{f}-t_{s}^{*}}{N}\sum_{i=1}^{N}l\left(\sigma_{i}(q)\right)$ (62)
$\leq v\left(t_{s}^{*}\right)-t_{s}^{*}l\left(\rho_{\min}\right)+\frac{\eta}{2}$
$\leq m^{*}+\left(t_{f}-t_{s}^{*}\right)l\left(\rho_{\min}\right)+\eta.$
Since $x_{N+1}(q)$ is feasible for the perturbated problem, we have $G(x_{N+1}(q))=0$ and $x_{N+1}(q)$ is feasible for Problem 7. Therefore, $m(x_{N+1}(q))\geq m^{*}$ and
$\frac{\sum_{i=1}^{N}l\left(\sigma_{i}(q)\right)}{N}\leq\frac{\eta}{t_{f}-t_{s}^{*}}+l\left(\rho_{\text{min}}\right).$ (63)
As $\sigma_{i}(q)\geq\rho_{\text{min}}$ for all $i$ and $l$ has positive derivative, we have
$l\left(\rho_{\text{min}}\right)\leq l\left(\sigma_{i}(q)\right)\leq l\left(\rho_{\text{min}}\right)+\frac{N\eta}{t_{f}-t_{s}^{*}}.$ (64)
Given that $\eta\leq\epsilon L_{l}(t_{f}-t_{s}^{*})/N$ and the derivative of $l$ is larger than $L_{l}$, we have
$\left|\sigma(q)-\rho_{\text{min}}\right|\leq\frac{N\eta}{L_{l}(t_{f}-t_{s}^{*})}\leq\epsilon,$ (65)
yielding the third result.
For the final result, since $l\left(\rho_{\text{min}}\right)\leq l\left(\sigma_{i}(q)\right)$, we have
$m\left(x_{N+1}(q)\right)\leq m^{*}+\eta.$ Furthermore, Theorem 22 indicates that
$|m(\tilde{x}_{N+1}(q))-m(x_{N+1}(q))|\leq NDL_{A}L_{M}\|B\|\|q\|,$
where $D$, $L_{A}$, and $L_{M}$ are defined in Theorem 22 and are constants when $t_{s}^{*}$ is fixed. Thus, we obtain
$m\left(\tilde{x}_{N+1}(q)\right)\leq m^{*}+\eta+NDL_{A}L_{M}\|B\|\|q\|.$ (66)
With a sufficiently small $\|q\|$, we have $m\left(\tilde{x}_{N+1}(q)\right)\leq m^{*}+\epsilon.$